### v1

# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

Abstract
The convergence rate of a Markov chain to its stationary distribution is typically assessed using the concept of total variation mixing time. However, this worst-case measure often yields pessimistic estimates and is challenging to infer from observations. In this paper, we advocate for the use of the average-mixing time as a more optimistic and demonstrably easier-to-estimate alternative. We further illustrate its applicability across a range of settings, from two-point to countable spaces, and discuss some practical implications.
Keywords— Ergodic Markov chain; average mixing time; countable state space; $\beta$-mixing.
1   Introduction
The total variation distance variant of the mixing time $t_{\mathsf{mix}}$ of a Markov chain $X_{1},\dots,X_{n}$ on a discrete state space $\mathcal{X}$, governed by a transition operator $P$, quantifies the time to convergence to its stationary distribution $\pi$. Two observations on the chain separated in time by about $t_{\mathsf{mix}}$ will be approximately independent and sampled from $\pi$. Estimates on the mixing time yield exponential concentration inequalities (Paulin, 2015), and in turn, generalization bounds for stable machine learning algorithms trained with Markovian data (Mohri and Rostamizadeh, 2010; Agarwal and Duchi, 2012; Oliveira et al., 2022), regret bounds for reinforcement learning algorithms (Ortner, 2020), and convergence diagnostics for Markov Chain Monte Carlo methods, which are widely used in mathematical physics (Metropolis et al., 1953; Hastings, 1970), computer science (Sinclair, 2012) and statistics (Chib and Greenberg, 1995).
However, $t_{\mathsf{mix}}$ is an intrinsically pessimistic notion, defined in terms of the worst-case initial distribution. Furthermore, $t_{\mathsf{mix}}$ is generally unknown a priori, and theoretical upper bounds, when available, are typically conservative. Finally, estimation of $t_{\mathsf{mix}}$ from a single trajectory of observations of length $n$ is known to be a statistically hard problem which requires at least
$n\in\Omega\left(\frac{1}{\pi_{\star}}+t_{\mathsf{mix}}\left|\mathcal{X}\right|\right),$ (1)
for estimation down to constant multiplicative error (Hsu et al., 2019, Theorem 3.1, Theorem 3.2), where $\left|\mathcal{X}\right|$ is the size of the space and $\pi_{\star}$ is the minimum stationary probability (refer to Section 1.4). The sample complexity lower bound in (1) is prohibitively large for vast state spaces, and even vacuous in the infinite case.
In this paper, we propose to address some of the above listed concerns about the worst-case $t_{\mathsf{mix}}$, by considering the average-mixing time $t_{\mathsf{mix}}^{\sharp}$ which recently came under the spotlight in (Espuny Díaz et al., 2022; Münch and Salez, 2022). Specifically, we will confirm that $t_{\mathsf{mix}}^{\sharp}$ is a more optimistic measure of convergence, in the sense where some chains can average-mix arbitrarily faster than they mix, even on a two-point space. Furthermore, we initiate the program of estimating $t_{\mathsf{mix}}^{\sharp}$ from a single trajectory and we demonstrate that it is statistically less costly than the problem pertaining to $t_{\mathsf{mix}}$. In fact, contrary to the worst-case problem, our analysis carries to a large subset of countable-state Markov chains. One should rightfully expect $t_{\mathsf{mix}}^{\sharp}$ to be less informative than $t_{\mathsf{mix}}$ since it quantifies a weaker notion of mixing. However, we confirm that knowledge of $t_{\mathsf{mix}}^{\sharp}$ remains of practical value in the sense where it can be used in known powerful decoupling techniques (Yu, 1994) and appear as a natural parameter controlling the large deviation rate for additive functionals. Along the way, we explore the connections between the average-mixing time and $\beta$-mixing, argued by Vidyasagar (2013); Mohri and Rostamizadeh (2010) to be “just the right” assumption for the analysis of machine learning algorithms with weakly-dependent data.
1.1   Related work
Average variants of the mixing time have recently made their apparition in the literature. For instance, Berestycki et al. (2018) analyzed random walks on the giant component of the Erdös–Rényi random graph $\mathcal{G}(n,p)$ with $p=\Theta(1/n)$, and showed that the mixing time when starting randomly from any vertex is the order $\mathcal{O}(\log n)$ while it was known to be of the order $\mathcal{O}(\log^{2}n)$ for the mixing time from a worst starting point. Following up, Espuny Díaz et al. (2022) extended the approach on a larger collection of graphs with small bottlenecks, and show applicability of their approach for the Newman–Watts small-world model and supercritically percolated expander graphs. In (Lovász and Kannan, 1999), the authors introduce the notion of a conductance averaged over subsets of states with various sizes, and demonstrate that it can lead to tighter convergence bounds for counting and volume computation problems. Finally, the average-mixing time $t^{\sharp}_{\mathsf{mix}}$, considered in (Münch and Salez, 2022) and the object of our study, is a measure of convergence of a Markov chain defined with respect to the chain being started from its stationary distribution (refer to Section 2), instead of the worst-case state.
In particular, we stress that the notion of average-mixing treated in this paper is different from the one discussed in Peres and Sousi (2015); Anderson et al. (2021), defined by taking an average over two successive steps on the chain.
1.2   Main contributions
We direct the reader to Section 1.4 for the notation.
1.2.1 Highlight the significance of the average-mixing time
Observing that the average-mixing time $t^{\sharp}_{\mathsf{mix}}$ recently analyzed by Münch and Salez (2022) corresponds to the stationary $\beta$-mixing time of the Markov chain,
$t^{\sharp}_{\mathsf{mix}}(\xi)=\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\beta(s)\leq\xi\right\},$
we briefly demonstrate how the average-mixing time (together with an a priori known order of convergence) can be used to quantitatively capture the convergence rate of a Markov chain in lieu of the worst-case mixing time. Specifically, we show (Lemma 2.3) that the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\delta)$ naturally appears in the sample complexity for controlling the deviation probability from 00 of a centered function $f\colon\mathcal{X}\to[-1,1]$. Namely, for $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ to hold with probability at least $1-\delta$, we show that under sub-exponential $\beta$-mixing of order $b$, it is sufficient to have a trajectory of
$n\geq\frac{\left(e/b\right)^{1/b}}{b\varepsilon^{2}}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\operatorname{polylog}\left(\frac{1}{\delta},\frac{1}{\varepsilon}\right),$
where $\operatorname{polylog}$ is a poly-logarithmic function of its parameters and $e$ is the Euler number. Similarly, in the polynomial $\beta$-mixing setting of order $b$, we show that for some natural constant $C\in\mathbb{R}_{+}$,
$n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
is sufficient.
1.2.2 Estimation of $\beta$-mixing coefficients
For a general time-discrete, time-homogeneous Markov chain over a countable state space, we establish a parametric rate of convergence for its individual $\beta$-coefficients from a single stationary trajectory of observations $X_{1},\dots,X_{n}$. Namely, for $s\in\mathbb{N}$, we construct an estimator $\widehat{\beta}(s)$ for $\beta(s)$, and we show (Theorem 3.1) that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
with
$\mathcal{B}_{p}^{(s)}\leq\left\|\beta^{(s)}\right\|_{1/p}^{1/p},\qquad\mathcal{J}_{p}^{(s)}=\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
and where for $x,x^{\prime}\in\mathcal{X}$, $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$. Here, $\mathcal{B}_{p}^{(s)}$ is a mixing parameter roughly controlled from above by the $\ell_{1/p}$ quasi-norm of the sequence of $\beta$-mixing coefficients of the $s$-skipped Markov chain, and $\mathcal{J}_{p}^{(s)}$ is an entropic quantity pertaining to the $s$-skipped Markov chain. Crucially, the parameter $p\in[1,\infty)$ is free and will allow us to balance the two factors in the upper bound.
1.2.3 Estimation of the average-mixing time
In the uniformly ergodic setting, for $\xi\in(0,1)$ and $\delta\in(0,1)$, we show (Theorem 3.4) that a single path length of
$n\geq\frac{t_{\mathsf{mix}}}{\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,3\xi/4}\right\}\operatorname{polylog}\left(\frac{1}{\xi},\frac{1}{\delta},t_{\mathsf{mix}}\right),$
where $t_{\mathsf{mix}}$ is the mixing time of $P$ and
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$
is sufficient in order to obtain an estimate $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ of the average-mixing time $t^{\sharp}_{\mathsf{mix}}(\xi)$ which satisfies
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(3\xi/2),t_{\mathsf{mix}}^{\sharp}(\xi/2)\right]\end{split}$
with probability at least $1-\delta$. Furthermore, for finite spaces, setting $\xi=1/4$ for simplicity, the parameter $p$ can be chosen so as to obtain an upper bound (Corollary 4.2) of
$n\geq Ct_{\mathsf{mix}}\log\max\left\{\left|\mathcal{X}\right|,\frac{t_{\mathsf{mix}}}{\delta}\right\},$
where $C\in\mathbb{R}_{+}$ is a universal constant. One can thus demonstrably estimate with high probability the average-mixing time sub-linearly in the number of states, answering a question posed by Luca Zanetti and John Sylvester (Zanetti and John, 2023). This discovery stands in stark contrast with the worst-case mixing time estimation problem, or the relaxation time estimation problem, for which the sample complexity necessarily depends inversely on the minimum stationary probability (Hsu et al., 2019; Wolfer and Kontorovich, 2024).
1.3   Outline
Section 2 introduces the average-mixing time, illustrates its significance and analyzes some of its properties. In Section 2.2, under geometric ergodicity and time-reversibility assumptions, we establish an upper bound of $t^{\sharp}_{\mathsf{mix}}$ in terms of the relaxation time of the chain, which is non-vacuous over countable state spaces and under suitable conditions on the stationary distribution. In Section 2.3 we recall some important connection between the average-mixing time and $\beta$-mixing properties of the chain.
Section 3 is dedicated to the problem of inferring the $\beta$-mixing coefficients and average-mixing time from a single trajectory of observations. In Section 3.1, we estimate the $\beta$-mixing coefficients of the chain; first individually and then as a sequence, with respect to the uniform norm. In Section 3.3, we specialize our analysis to the uniformly ergodic setting, and in Section 3.4 we deduce an upper bound for the problem of estimating the average-mixing time from a single trajectory under uniform ergodicity.
In Section 4, we demonstrate that our contribution holds implications at various scales of state spaces and provide more explicit rates under additional structural assumptions on the underlying chain.
Technical proofs have been deferred to Section 5 for readability.
Finally, Section 6 introduces a class of Markov chain which illustrates some of our results.
1.4   Notation and setting
We let $\mathcal{X}$ be a countably infinite space of symbols equipped (for simplicity) with a total order. Real functions on $\mathcal{X}$ are written as infinite row vectors and for $x\in\mathcal{X}$, $e_{x}$ is the function defined by $x^{\prime}\in\mathcal{X}$, $e_{x}(x^{\prime})=\delta\left[x=x^{\prime}\right]$. For $p\in(0,\infty)$, and a function $v\in\mathbb{R}^{\mathcal{X}}$,
$\left\|v\right\|_{p}\doteq\left(\sum_{x\in\mathcal{X}}\left|v(x)\right|^{p}\right)^{1/p},$
for convenience, we define
$\left\|v\right\|_{0}\doteq\left|\left\{x\in\mathcal{X}\colon\left|v(x)\right|>0\right\}\right|,$
and for any $p\in[0,\infty)$,
$\ell_{p}\doteq\left\{s\in\mathbb{R}^{\mathcal{X}}\colon\left\|s\right\|_{p}<\infty\right\}.$
The set of all probability distributions over $\mathcal{X}$, written $\mathcal{P}(\mathcal{X})$, corresponds to non-negative functions summing to unity. A time-discrete, time-homogeneous Markov chain
$X_{1},X_{2},X_{3}\dots$
on $\mathcal{X}$ is defined by a transition operator represented by an infinite stochastic matrix $P$
$P(x,x^{\prime})=\mathbb{P}\left(X_{t+1}=x^{\prime}|X_{t}=x\right).$
We write $\mathcal{W}(\mathcal{X})$ for the set of all infinite stochastic matrices over $\mathcal{X}$. For a finite trajectory of length $n\in\mathbb{N}$, started from an initial distribution $\mu\in\mathcal{P}(\mathcal{X})$ and evolving according to $P\in\mathcal{W}(\mathcal{X})$, the probability of observing $(x_{1},x_{2},\dots,x_{n})\in\mathcal{X}^{n}$ can be factored as
$\mathbb{P}_{\mu}\left(X_{1}=x_{1},\dots,X_{n}=x_{n}\right)=\mu(x_{1})\prod_{t=1}^{n-1}P(x_{t},x_{t+1}).$
By abuse of notation, matrix multiplication will retain its meaning in the countable setting. Namely, for $\mu\in\mathcal{P}(\mathcal{X})$ and $P\in\mathcal{W}(\mathcal{X})$, we will write,
$\mu P=\sum_{x,x^{\prime}\in\mathcal{X}}\mu(x)P(x,x^{\prime})e_{x^{\prime}}.$
Extensions of notation to $Pf^{\intercal}$ with $f\in\ell_{1}$, or for $P^{t}$ are similarly defined. We say that $P$ is ergodic when $P$ is irreducible, aperiodic and positive-recurrent (Levin et al., 2009). In that case, there exists a unique stationary distribution $\pi$ verifying $\pi P=\pi$, and we can define the distance to stationarity
$d(t)\doteq\sup_{x\in\mathcal{X}}\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
where $\left\|\cdot\right\|_{\mathsf{TV}}$ is the total variation distance defined for $\mu,\nu\in\mathcal{P}(\mathcal{X})$ by $\left\|\mu-\nu\right\|_{\mathsf{TV}}\doteq\frac{1}{2}\left\|\mu-\nu\right\|_{1}$. We say that $P$ is uniformly ergodic whenever, there exist $M>0$ and $\rho\in(0,1)$ such that for any $t\in\mathbb{N}$,
$d(t)\leq M\rho^{t}.$
We write $\mathcal{W}^{\star}(\mathcal{X})$ the set of uniformly ergodic Markov chains. We recall that the (worst-case) mixing time is defined as
$\begin{split}t_{\mathsf{mix}}(\xi)\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d(t)\leq\xi\right\}.\end{split}$
It is furthermore customary to define
$\qquad t_{\mathsf{mix}}\doteq t_{\mathsf{mix}}(1/4),$ (2)
as a consequence of the following property (Levin et al., 2009, Lemma 4.11)
$t_{\mathsf{mix}}(\xi)\leq\left\lceil\log_{2}\frac{1}{\xi}\right\rceil t_{\mathsf{mix}}.$ (3)
Elements related directly to the notion of average-mixing time will be introduced in the next section for clarity of the exposition.
2   The average-mixing time
In this section, we first recall the definition of the average-mixing time, and subsequently explore some of its basic properties. For $P\in\mathcal{W}(\mathcal{X})$, the average distance to stationarity of the chain induced from $P$ at time $t\in\mathbb{N}$ is expressed as
$d^{\sharp}(t)\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{t}-\pi\right\|_{\mathsf{TV}},$
where $\pi$ is the stationary distribution of $P$. The above should be understood as the total variance distance of the chain to stationarity, averaged with respect to $\pi$ over the starting state. The limit $\lim_{t\to\infty}d^{\sharp}(t)=0$ always exists [^1] This follows from a connection between average-mixing time and $\beta$-mixing discussed at Section 2.3, thus for $\xi\in(0,1)$, the average-mixing time of $P$, defined (Münch and Salez, 2022) as
$\begin{split}t^{\sharp}_{\mathsf{mix}}(\xi)&\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{d^{\sharp}(t)\leq\xi\right\},\end{split}$
always verifies $t^{\sharp}_{\mathsf{mix}}(\xi)<\infty$.
Remark 2.1.
Similar to $t_{\mathsf{mix}}$ in (2), it is tempting to introduce $t_{\mathsf{mix}}^{\sharp}\doteq t_{\mathsf{mix}}^{\sharp}(1/4)$. However, since $d^{\sharp}$ does not enjoy the same sub-multiplicativity properties as $d$ (3) (refer to (Levin et al., 2009, Lemma 4.11)), it is generally not enough to consider a fixed value of $\xi\in(1/2)$, and in this manuscript we will keep $\xi$ to be a user-fixed value.
2.1   Average-mixing versus worst-case mixing
Following their respective definitions, it holds that $d^{\sharp}(t)\leq d(t)$. Thus, for any fixed $\xi\in(0,1)$, average-mixing is always faster than mixing. The equality $t_{\mathsf{mix}}=t^{\sharp}_{\mathsf{mix}}$ is notably achieved for transitive chains (Münch and Salez, 2022, Remark 1), which in particular include random walks on groups [^2] A chain is called transitive when for any $x_{0},x_{0}^{\prime}\in\mathcal{X}$, there exists a bijection $\phi\colon\mathcal{X}\to\mathcal{X}$, with $\phi(x_{0})=x_{0}^{\prime}$ and for all $x,x^{\prime}\in\mathcal{X},P(\phi(x),\phi(x^{\prime}))=P(x,x^{\prime})$.. In general however, the mixing time can be dramatically larger than the average-mixing time, even when the state space is small, as formalized below.
Proposition 2.1.
Let $M\in\mathbb{R}_{+}$ be arbitrarily large. There exists a transition operator $P$ such that for any $\xi\in(0,1)$, it holds that
$t_{\mathsf{mix}}(\xi)>Mt_{\mathsf{mix}}^{\sharp}(\xi).$
Proof.
Lemma 4.6 establishes this fact in the case where $\mathcal{X}=\left\{0,1\right\}$. ∎
2.2   Spectral methods under reversibility and geometric ergodicity
In this section, we endow $\mathbb{R}^{\mathcal{X}}$ with the inner product $\langle f,g\rangle_{\pi}\doteq\sum_{x\in\mathcal{X}}f(x)g(x)\pi(x)$, where $\pi$ is a positive distribution. We regard $P$ as a linear operator on the resulting Hilbert space $\ell_{2}(\pi)\doteq(\mathbb{R}^{\mathcal{X}},\langle\cdot,\cdot\rangle_{\pi})$ and further assume $P$ to be geometrically ergodic. Recall that $P\in\mathcal{W}^{\star}(\mathcal{X})$ is called $\pi$-reversible when the following detailed-balance equation holds,
$\pi(x)P(x,x^{\prime})=\pi(x^{\prime})P(x^{\prime},x),\qquad\forall x,x^{\prime}\in\mathcal{X},$
which is equivalent to stating that $P$ is self-adjoint in $\ell_{2}(\pi)$. Under reversibility, the spectrum of $P$, denoted $\operatorname{Spec}(P)$, is contained in the real line. The absolute spectral gap of $P$ is then defined as
$\gamma_{\star}\doteq 1-\sup\left\{\left|\lambda\right|\colon\lambda\in\operatorname{Spec}(P),\lambda\neq 1\right\}.$
Geometric ergodicity and existence of a positive spectral gap are known to be equivalent (Kontoyiannis and Meyn, 2012). We first establish that, at least under mild assumptions on $P$, the average-mixing time can be controlled by the relaxation time $t_{\mathsf{rel}}\doteq 1/\gamma_{\star}$, even over countably infinite spaces.
Lemma 2.1.
Let $P\in\mathcal{W}^{\star}(\mathcal{X})$ be $\pi$-stationary, geometrically ergodic, reversible and let $\xi\in(0,1)$. When $P$ regarded as the linear operator $P\colon\ell_{2}(\pi)\to\ell_{2}(\pi)$ is compact [^3] For instance, being trace-class is a sufficient condition for $P$ to be compact., it holds that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{2\xi},$
where $t_{\mathsf{rel}}$ is the relaxation time of $P$.
The half quasi-norm of the stationary distribution $\|\pi\|_{1/2}$, making its first appearance in our study, is a natural measure of complexity of the stationary distribution $\pi$, directly related to its Rényi entropy of order $1/2$. Note that $\|\pi\|_{1/2}$ could be infinite, in which case Lemma 2.1 becomes vacuous. However, Lemma 2.1 compares favorably with the bound of Münch and Salez (2022, Theorem 5, Corollary 2), which is inversely proportional to the minimum (non-zero) transition probability $P_{\star}\doteq\min_{x,x^{\prime}\in\mathcal{X}}P(x,x^{\prime})$. Additionally, Lemma 2.1 can be instructively compared with Levin et al. (2009, Theorem 12.4) for the mixing time $t_{\mathsf{mix}}(\xi)$,
$t_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{1}{\xi\pi_{\star}},$
where $\pi_{\star}\doteq\min_{x\in\mathcal{X}}\pi(x)$, which is only informative [^4] We note in passing that the bound at Levin et al. (2009, Theorem 12.4) could be improved as $t_{\mathsf{mix}}(\xi)\leq\frac{1}{2}t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{\xi\pi_{\star}}$. when $\left|\mathcal{X}\right|<\infty$.
2.3   Connection with $\beta$-mixing
Recall that for a process $\{X_{t}\}_{t\in\mathbb{Z}}$ on $\mathcal{X}$, for any $s\in\mathbb{Z}$, the $\beta$-mixing coefficient (Bradley, 2005; Doukhan, 2012) is defined by
$\beta(s)\doteq\sup_{r\in\mathbb{N}}\beta\left(\sigma\left(\left\{X_{t}\colon t\leq r\right\}\right),\sigma\left(\left\{X_{t}\colon t\geq r+s\right\}\right)\right),$
where $\sigma(\{X_{t}\}_{t})$ is the $\sigma$-field generated by the random variables $\{X_{t}\}_{t}$, and for two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$,
$\beta(\mathcal{A},\mathcal{B})\doteq\sup\frac{1}{2}\sum_{i=1}^{I}\sum_{j=1}^{J}\left|\mathbb{P}(A_{i}\cap B_{i})-\mathbb{P}(A_{i})\mathbb{P}(B_{j})\right|,$
where the supremum is taken over all pairs of finite partitions $\{A_{1},\dots,A_{I}\}$ and $\{B_{1},\dots,B_{J}\}$ of $\mathcal{X}$ such that for any $1\leq i\leq I$ and any $1\leq j\leq J$, $A_{i}\in\mathcal{A}$ and $B_{j}\in\mathcal{B}$ (Bradley, 2005, (7)). The random process $\{X_{t}\}_{t\in\mathbb{Z}}$ is called $\beta$-mixing, or absolutely regular, when
$\lim_{s\to\infty}\beta(s)=0.$
Recall that a stationary, countable-state Markov chain $X$ is $\beta$-mixing if and only if it is irreducible and aperiodic (Bradley, 2005, Theorem 3.2).
Lemma 2.2 (Davydov (1974, Proposition 1)).
We let $\{X_{t}\}_{t\in\mathbb{N}}$ be a countable-state Markov chain with transition operator $P$. Writing $\mu^{(0)}$ for the initial distribution, and for $t>1$, $\mu^{(t)}=\mu^{(0)}P^{t}$ for the marginal distribution after time step $t$, it holds that
$\beta(s)=\sup_{t\in\mathbb{N}}\sum_{x\in\mathcal{X}}\mu^{(t)}(x)\left\|e_{x}P^{s}-\mu^{(s+t)}\right\|_{\mathsf{TV}}.\$
It follows from the above lemma that when the chain is stationary, $\mu^{(0)}=\pi=\mu^{(t)}$ for any $t\in\mathbb{N}$, and
$\begin{split}\beta(s)=\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{\mathsf{TV}}=d^{\sharp}(s),\\ \end{split}$
thus the average-mixing time is essentially the “stationary $\beta$-mixing time” of the chain555 Similarly, the worst-case mixing time can be connected to the notion of $\phi$-mixing (Davydov, 1974). . Note that with the convention $P^{0}=I$, we obtain
$\beta(0)=1-\left\|\pi\right\|_{2}^{2},$
which depends only on the stationary properties of the chain. When there exists $\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{1}/s^{b}$, we say that the process $\beta$-mixes algebraically or polynomially. When there exists $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ we say that the chain $\beta$-mixes sub-exponentially fast when $b\neq 1$, and exponentially fast when $b=1$. In Section 6, we will introduce a class of Markov transition operators, whose flavor of mixing (algebraic, sub-exponential, exponential) can be fully controlled. Notably (refer to the proof of Theorem 3.2), when $P$ is uniformly ergodic with mixing time $t_{\mathsf{mix}}$, it holds that
$\beta(s)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$ (4)
hence the process is exponentially $\beta$-mixing with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. However, nothing a priori precludes the $\beta$-mixing rate from being superior to (4). The connection between the average-mixing time and $\beta$-mixing provides additional motivation for the problem of estimating the $\beta$-mixing coefficients, whose knowledge is known to lead to powerful decoupling techniques (Yu, 1994). The problem of inferring mixing coefficients from a single trajectory of observations has recently received attention from the community. In the most recent work (Khaleghi and Lugosi, 2021), consistent estimators for $\beta$- and $\alpha$-mixing coefficients were proposed. In Section 3, we will provide more explicit and finer finite-sample bounds for the estimation problem in the specialized Markovian setting. A regime of particular interest to us is when the chain $\beta$-mixes much faster than $\Theta(t_{\mathsf{mix}})$. Henceforth in this manuscript, when they coincide under stationarity, we may use the notations $\beta$ and $d^{\sharp}$ interchangeably.
2.4   Bounding the deviation of additive functionals evaluated on Markov chains
We now illustrate how the average-mixing time, through its connection to $\beta$-mixing naturally appears as a quantity of interest when bounding deviations of functions evaluated on a trajectory sampled from a Markov chain. More formally, let $X_{1},\dots,X_{n}$ be a Markov chain over a countable space $\mathcal{X}$, and whose dynamics are governed by $P\in\mathcal{W}(\mathcal{X})$ with stationary distribution $\pi$. For simplicity, we consider a real bounded function $f$ on $\mathcal{X}$, which is centered in the sense where its stationary expected value vanishes $\mathbb{E}_{\pi}f=0$. The ergodic theorem states that the empirical mean taken on an infinite trajectory will similarly vanish,
$\lim_{n\to\infty}\frac{1}{n}\sum_{t=1}^{n}f(X_{t})=0.$
In practice however, we are more likely to be interested in controlling the deviation of the sample mean from its expected value in a non-asymptotic setting.
Lemma 2.3 (Bounding deviation of additive functionals evaluated on Markov chains).
Let $f\colon\mathcal{X}\to[-1,1]$ with $\mathbb{E}_{\pi}f=0$, let $\varepsilon,\delta\in(0,1)$, and let $t^{\sharp}_{\mathsf{mix}}\colon(0,1)\to\mathbb{N}$. For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, and average-mixing time bounded from above by $t^{\sharp}_{\mathsf{mix}}$, the following statements hold.
  1. 1.
    (Sub-exponential) If there exist $\beta_{0},\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{0}\geq 1$, $b\in(0,1]$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
    $n\geq\frac{C}{\varepsilon^{2}}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right)\log\left(\frac{1}{\delta}\right),$
    with
    $\xi(\varepsilon,\delta)\doteq\frac{\delta\varepsilon^{2}}{16\log\left(\frac{4}{\delta}\right)},$
    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$. The above upper on the path length can be further streamlined as
    $n\geq\frac{C}{\varepsilon^{2}}\frac{1}{b}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\delta/e\right)\log\left(\frac{1}{\delta}\right)\log\left(\frac{\log(1/\delta)}{\varepsilon}\right),$
    for some universal constant $C\in\mathbb{R}_{+}$.
  2. 2.
    (Polynomial) If there exist $\beta_{1},b\in\mathbb{R}_{+}$ with $\beta_{1}\geq 1$, such that for any $s\in\mathbb{N}$, $\beta(s)\leq\frac{\beta_{1}}{s^{b}}$, then there exists a universal constant $C\in\mathbb{R}_{+}$ such that for
    $n\geq C\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(2t^{\sharp}_{\mathsf{mix}}\left(\delta\right)\right)^{\frac{1}{b}},$
    it holds that $\frac{1}{n}\sum_{t=1}^{n}f(X_{t})\leq\varepsilon$ with probability at least $1-\delta$.
Proof sketch.
We adopt the traditional method as described in Yu (1994). This approach combines a blocking technique, credited to Bernstein (1927) in the context of $m$-dependent processes, together with an application of Hoeffding’s inequality to control the deviation probability of a sum of independent observations. We further optimize for the block size and reformulate the results to obtain an expression involving the average-mixing time under various rates of $\beta$-mixing. Refer to Section 5.1 for the proof. ∎
As a consequence of Lemma 2.3 , in both the sub-exponential and polynomial settings, provided an assumption on the convergence order $b$, estimates on the average-mixing time to proximity $\Theta(\delta)$ yield corresponding deviation bounds. In the uniformly ergodic setting, corresponding to the assumption of the exponential setting ($b=1$), we obtain a deviation inequality involving the average-mixing time in lieu of the mixing time (Chung et al., 2012, Theorem 3.1). The lemma generalizes to non-centered functions, lower deviations and arbitrarily bounded functions. It is also noteworthy that in the sub-exponential setting, the convergence order $b$ solely affects the constant in the sample complexity. Our formulation above serves as additional motivation for analyzing and inferring the average-mixing time.
3   Estimation from a single trajectory
In this section, we analyze the estimation of $\beta$-mixing coefficients and the average-mixing time from a single trajectory of observations.
3.1   Estimation of $\beta$-mixing coefficients under general ergodicity condition
In this section we analyse plug-in estimators for $\beta$-mixing coefficients. We recall that in this general setting, the mixing time can be infinite, and the $\beta$-mixing coefficients can decay at arbitrary slow rate. We will focus on exponential, sub-exponential and polynomial rates of convergence. Our analysis relies on stationary skipped trajectories, where for $s\in\mathbb{N},s<n$, we denote
$\begin{split}X^{(s)}\doteq X_{1},X_{1+s},X_{1+2s},\dots,X_{1+\lfloor(n-1)/s\rfloor s},\end{split}$
with transition operator $P^{s}$ and initial distribution $\pi$. It will be convenient to define the following counting random variables,
$\begin{split}N_{x}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x\right],\\ N_{xx^{\prime}}^{(s)}&\doteq\sum_{t=1}^{\lfloor(n-1)/s\rfloor}\delta\left[X_{1+s(t-1)}=x,X_{1+st}=x^{\prime}\right].\end{split}$
We use the shortcut notation $N_{x}=N_{x}^{(1)}$ and $N_{xx^{\prime}}=N_{xx^{\prime}}^{(1)}$ to respectively denote number of visits to state $x$ and transitions from $x$ to $x^{\prime}$ on the original chain. We define the following estimator $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ for $\beta(s)$,
$\widehat{\beta}(s)=\frac{1}{2\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|,$ (5)
and by convention, we set $\widehat{\beta}(s)=0$ for $s\geq n$.
Theorem 3.1.
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$, and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$, it holds that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{\mathcal{B}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\frac{3}{2}\mathcal{J}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}}\right)\leq 4\sqrt{\frac{\mathcal{B}_{p}^{(s)}\mathcal{J}_{p}^{(s)}}{\lfloor(n-1)/s\rfloor}},\end{split}$
where for $x,x^{\prime}\in\mathcal{X}$, we wrote $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$, and
$\mathcal{B}_{p}^{(s)}\doteq\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p},\qquad\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}.$
We observe that the two key quantities $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$ control the estimation rate. The former essentially captures $\beta$-mixing properties of the process, while the latter is tightly connected to entropic properties of the distribution of (skipped) stationary pairs. In the following Corollary 3.1, we readily provide more explicit upper bounds on $\mathcal{B}_{p}^{(s)}$ under additional assumptions on the mixing rate. The analysis of $\mathcal{J}_{p}^{(s)}$ is deferred to Section 4.
Corollary 3.1.
Let $s,n\in\mathbb{N}$ with $s<n$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$ with $\mathcal{J}_{p}^{(s)}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}<\infty$, the following statements hold.
  1. 1.
    (Exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$ and $t\in\mathbb{N}\cup\left\{0\right\}$,
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}\left(\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}\right)\mathcal{J}_{p}^{(s)}}.\\ \end{split}$
  2. 2.
    (Sub-exponential) If $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$, with $\beta_{0},\beta_{1}\in\mathbb{R}_{+}$, $b\in(0,1)$, and $s\in\mathbb{N}\cup\left\{0\right\}$.
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\exp(-\beta_{1}s^{b}/p)+\frac{\Gamma\left(\frac{1}{b},s^{b}\beta_{1}/p\right)}{bs(\beta_{1}/p)^{1/b}}\right)\mathcal{J}_{p}^{(s)}},\end{split}$
[…]
  3. 3.
    (Polynomial) If $\beta(s)\leq\beta_{1}/s^{b}$, with $\beta_{1}\in\mathbb{R}_{+}$, $b\in(1,\infty)$, $b/p\in(1,\infty)$, $s\in\mathbb{N}$ and $\beta(0)=\beta_{0}\in\mathbb{R}_{+}$,
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{\beta_{0}^{1/p}+\zeta(b/p)\frac{\beta_{1}^{1/p}}{s^{b/p}}}{\lfloor(n-1)/s\rfloor}\mathcal{J}_{p}^{(s)}},\end{split}$
    where $\zeta$ is the Riemann zeta function defined by $\zeta(r)=\sum_{t=1}^{\infty}t^{-r}$.
3.2   Comparison with prior results
In McDonald et al. (2015, Theorem 4), the authors establish the nearly parametric convergence rate for continuously valued Markov processes
$\mathbb{E}{\left|\widehat{\beta}(s)-\beta(s)\right|}\leq\mathcal{O}\left(\sqrt{\frac{W_{0}(n)}{n}}\right),$
where $W_{0}$ is the Lambert function (Corless et al., 1996). They also posit that their findings may be considered conservative when applied to discrete-valued processes. Theorem 3.1 validates their remark by recovering the superior parametric rate $\mathcal{O}(\sqrt{1/n})$ under convergence of $\mathcal{B}_{p}^{(s)}$ and $\mathcal{J}_{p}^{(s)}$. Our method shares similarities with Khaleghi and Lugosi (2021), as both employ Rio’s covariance inequality Rio (1999, Corollary 1.1). However, compared to both Khaleghi and Lugosi (2021) and McDonald et al. (2015, Theorem 4), we distinguish ourselves by focusing on the discrete space setting and deriving explicit rates in terms of fine-grained properties of the transition operator.
3.3   Estimation of $\beta$-mixing coefficients under uniform ergodicity
Under uniform ergodicity, the mixing time is finite, and the $\beta$-mixing coefficients decays exponentially. In this setting, we show how to obtain a sample complexity upper bound for high-probability estimates on the invividual $\beta$-mixing coefficients.
Theorem 3.2.
Let $s,n\in\mathbb{N}$ with $n>2s+1$, let $p\in\mathbb{R}_{+},p\geq 1$ and let $\widehat{\beta}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined in (5). For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$,
  1. $(i)$
    $\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 8\sqrt{\frac{1}{n-1}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
    where $t_{\mathsf{mix}}$ is the mixing time of $P$.
  2. $(ii)$
    Let $\varepsilon,\delta\in(0,1)$. For
    $n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{2}{\delta}\right\},$
    with probability at least $1-\delta$, it holds that
    $\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon.$
Using a truncated estimator with a cut-off determined from a uniform bound on the complexity terms $\|Q^{(s)}\|_{1-1/p}^{1-1/p}$, we can control the sequence of $\beta$-mixing coefficients with respect to the uniform norm modulo an additional logarithmic factor in the sample complexity.
Theorem 3.3.
Let $\varepsilon\in(0,1/4],\delta\in(0,1),p\in\mathbb{R}_{+},p\geq 1$ and $\mathcal{J}_{p}\in\mathbb{R}_{+}$. Let $n,S\in\mathbb{N}$, and for $s\in\mathbb{N}$, let $\widehat{\beta}_{[S]}(s)\colon\mathcal{X}^{n}\to[0,1]$ be the estimator defined as
$\widehat{\beta}_{[S]}(s)=\begin{cases}\widehat{\beta}(s)&\text{ for }s\leq S\\ 0&\text{ otherwise,}\end{cases}$
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
$\sup_{s\in\mathbb{N}}\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\leq\mathcal{J}_{p},$
when
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\varepsilon}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that,
$\left\|\widehat{\beta}_{[S(n,\varepsilon,\delta,p,\mathcal{J}_{p})]}-\beta\right\|_{\infty}\leq\varepsilon,$
where the cut-off is set as
$S(n,\varepsilon,\delta,p,\mathcal{J}_{p})\doteq\frac{(n-1)\varepsilon^{2}}{2^{9}\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}}.$
Remark 3.1.
We note but do not pursue the fact that a more delicate bound on the sample complexity in Theorem 3.3 could be obtained, replacing a logarithm with a Lambert $W$ function.
In Section 4, we show how to obtain bounds on the entropy term in several natural settings.
3.4   Estimation of the average-mixing time
We move on to the estimation of the average-mixing time. Since the average distance to stationarity $d^{\sharp}=\beta$ does not generally enjoy sub-multiplicativity (it does not hold that $t^{\sharp}_{\mathsf{mix}}(\xi)(1/4)$ can be bounded by $t^{\sharp}_{\mathsf{mix}}$ and a polynomial of $\log\frac{1}{\xi}$), we formalize desired guarantees somewhat differently from Hsu et al. (2019); Wolfer and Kontorovich (2024). Indeed, for the relaxation time $t_{\mathsf{rel}}$, we were seeking an estimator $\widehat{t_{\mathsf{rel}}}$ to multiplicative error $\varepsilon$. Namely, the estimators in Hsu et al. (2019); Wolfer and Kontorovich (2024) guarantee that with high probability,
$\left|\frac{\widehat{t_{\mathsf{rel}}}}{t_{\mathsf{rel}}}-1\right|\leq\varepsilon.$
Here, we keep the distance to stationarity as a free parameter, and construct an estimator for the average-mixing time for proximity parameter $\xi$ up to multiplicative accuracy $\varepsilon$. Define for $\xi\in(0,1)$,
$\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\doteq\operatorname*{arg\,min}_{s\in\mathbb{N}}\left\{\widehat{\beta}(s)\leq\xi\right\},$ (6)
where $\widehat{\beta}(s)$ is the estimator for $\beta(s)$ defined in (5).
Theorem 3.4.
Let $\varepsilon\in(0,1/4),\delta,\xi\in(0,1),p\in\mathbb{R}_{+}$ and $\mathcal{J}_{p,\xi}\in\mathbb{R}_{+}$. Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6). For any uniformly ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}^{\star}(\mathcal{X})$ such that,
$\sup\left\{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\}\leq\mathcal{J}_{p,\xi},$
when
$n\geq 1+\frac{2^{8}t_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+\frac{p}{\log 2}\right)\mathcal{J}_{p,\frac{3\xi}{4}},9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
3.5   Extension to non-stationary Markov chains
Although results in this section are stated assuming a stationary start, it is possible to generalize them to non-stationary chains as follows. Let $\mu\in\mathcal{P}(\mathcal{X})$ absolutely continuous with respect to $\pi$. Inspired by the proof of Paulin (2015, Proposition 3.10), an application of Hölder’s inequality with $q,r\geq 1$ with $11/q+1/r=$ 1 yields,
$\mathbb{E}_{\mu}f(X_{1},\dots,X_{n})\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\left\|f(X_{1},\dots,X_{n})\right\|_{\pi,r},$
and in particular, for any $\varepsilon>0$, we can bound the probability of an $\varepsilon$-deviation assuming a non-stationary start as
$\mathbb{P}_{\mu}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)\leq\left\|\mu(X)/\pi(X)\right\|_{\pi,q}\mathbb{P}_{\pi}\left(f(X_{1},\dots,X_{n})>\varepsilon\right)^{1/r}.$
Setting $q=r=2$ recovers Paulin (2015, Proposition 3.10). The reader is invited to consult (Paulin, 2015, Section 3.3) for additional methods involving “burning” a prefix of the observed trajectory.
3.6   Comparison with an alternative method
We point out the fact that for the special case $p=2$, by using the techniques in Paulin (2015), one can obtain an upper bound of the order (ignoring logarithmic factors) of
$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\mathcal{J}_{2,\xi}$
for the problem of estimating the average-mixing time in uniformly ergodic Markov chains. Indeed, the mixing time of the $s$-skipped pairs of observations is upper bounded by $t_{\mathsf{mix}}^{(s)}+s\leq\lceil t_{\mathsf{mix}}/s\rceil+s$. From Paulin (2015, Theorem 3.2, Proposition  3.4), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor\lceil t_{\mathsf{mix}}/s\rceil\pi(x),\\ \operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]&\leq 2\lfloor(n-1)/s\rfloor(\lceil t_{\mathsf{mix}}/s\rceil+s)\pi(x)P^{s}(x,x^{\prime}),\\ \end{split}$
which yields
$\begin{split}\mathbb{E}\left|\widehat{\beta}(s)-\beta(s)\right|\leq&\sqrt{2\frac{\lceil t_{\mathsf{mix}}/s\rceil+s}{\lfloor(n-1)/s\rfloor}}\left(\left\|Q^{(s)}\right\|_{1/2}^{1/2}+2\left\|\pi\right\|_{1/2}^{1/2}\right).\\ \end{split}$
Our method is superior in that it introduces a trade-off parameter $p$ to obtain an upper bound of order
$\frac{t_{\mathsf{mix}}}{\varepsilon^{2}\xi^{2}}\inf_{p\geq 1}\left\{p\mathcal{J}_{p,\xi}\right\}.$
This enables the use of more versatile bounds on the entropic term, which we explore in Section 4.
4   Implications at various scales of state space
In infinitely denumerable state spaces, recall that for any $s\in\mathbb{N}$, the well-definition of the estimation rate for the $\beta(s)$-mixing coefficient (Theorem 3.1) hinges on the finiteness of the following quantity
$\mathcal{J}^{(s)}_{p}\doteq\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
where $p\in[1,\infty)$ is a user-fixed parameter. It is instructive to observe that on one hand
$\lim_{p\to\infty}\mathcal{J}_{p}^{(s)}=1,$
while on the other hand, $\mathcal{B}_{p}^{(s)}$ increases with $p$,
$\lim_{p\to\infty}\mathcal{B}_{p}^{(s)}\asymp\frac{n}{s},$
leading to a trade-off in the choice of the parameter in Theorem 3.1. In particular, provided a control which holds uniformly over $s$,
$\mathcal{J}_{p,\xi}\doteq\sup\left\{\mathcal{J}^{(s)}_{p}\colon s\in\mathbb{N},s\leq t^{\sharp}_{\mathsf{mix}}(\xi)\right\},$ (7)
the free parameter $p$ can be appropriately selected in Theorem 3.4, in order to optimize the upper bound on the sample complexity of estimating the average-mixing time. It is also noteworthy that since the $\ell_{1-1/p}$ quasi-norm and the $(1-1/p)$th order Rényi entropy $H_{1-1/p}$ of a distribution $\mu\in\mathcal{P}(\mathcal{X})$ are related by the expression
$p\log\left\|\mu\right\|_{1-1/p}^{1-1/p}=H_{1-1/p}(\mu),$
the quantity $\mathcal{J}_{p,\xi}$ pertains to a specific entropic property inherent in the Markov process. Unfortunately, this term is generally not easily amenable to analysis; hence, our focus lies in establishing bounds for its evaluation. Under ergodicity of $P$, the pair-wise limit
$\lim_{s\to\infty}Q^{(s)}(x,x^{\prime})=\pi(x)\pi(x^{\prime}),\qquad\forall x,x^{\prime}\in\mathcal{X},$
hints at a heuristic lower bound in $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ on $\mathcal{J}_{p,\xi}$. Using the stationary regime as a baseline, it seems natural to examine chains where the entropic term does not undergo relative explosive growth in the transient regime. For instance, for some multiplicative factor $K\geq 1$, an upper bound
$\mathcal{J}_{p}^{(s)}\leq K\left\|\pi\right\|^{2(1-1/p)}_{1-1/p},$ (8)
with a controlled $K$ readily simplifies the sample complexities appearing in Section 3. In the present section, we proceed to discuss bounds on the entropic quantities $\mathcal{J}_{p}^{(s)}$ and $\mathcal{J}_{p,\xi}$ in terms of state-wise conditional entropy, structural properties of the connection graph of the Markov chain, or the alphabet size. We conclude with a discussion on the binary state space, where the dominant complexity stems from the mixing bottlenecks, revealing that our results hold implications at every scale.
4.1   Ergodic properties
In this subsection, we derive various controls of $\mathcal{J}_{p}^{(s)}$ based on ergodic properties of the chain. Specifically, we introduce three conditions. Condition 1 is the classical notion of $V$-geometric ergodicity, which is known to be weaker than uniform ergodicity. Condition 2 imposes a stronger criterion, under which an ergodic chain is necessarily uniformly ergodic. Both Conditions 1 and 2 allow to derive bounds on $\mathcal{J}_{p}^{(s)}$ that are uniform over $s$. However, these conditions are quite stringent. We introduce a natural extension of these two conditions, Condition 3, which also allows to control $\mathcal{J}_{p}^{(s)}$. However, under Condition 3, this bound might be growing (sub-linearly) with $s$. We summarize the relations between these conditions in Figure 1.
{tikzcd}
Figure 1: Relation between the various ergodic conditions in Subsection 4.1. In this diagram, $\Rightarrow$ is to be read as ”implies under the assumptions of the result referred to”.
Condition 1 ($V$-geometric ergodicity (Douc et al., 2018, Section 18)).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is $V$-geometrically ergodic for some function $V\colon\mathcal{X}\to[1,\infty)$, when there exist constants $\theta\in(0,1)$ and $M\in\mathbb{R}_{+}$ such that for any $x\in\mathcal{X}$,
$\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s},$
where for any $v\in\mathbb{R}^{\mathcal{X}}$, we wrote
$\left\|\mu\right\|_{V}\doteq\sup_{x\in\mathcal{X}}V(x)v(x).$
Condition 2 (Existence of a dominating distribution).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that there exists a universal constant $C\in\mathbb{R}_{+}$ and a probability measure $\nu\in\mathcal{P}(\mathcal{X})$, such that for all $x^{\prime}\in\mathcal{X}$,
$\sup_{x\in\mathcal{X}}P(x,x^{\prime})\leq C\nu(x^{\prime}).$
Condition 3 (Weak $V$-geometric ergodicity).
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. We say that $P$ is weakly $V$-geometrically ergodic when there exists functions
$V\colon\mathcal{X}\to[1,\infty),\qquad D\colon\mathcal{X}\to[1,\infty),$
and a constant $b\geq 0$ such that for any $s\in\mathbb{N}$,
$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P^{s}(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)+b(s-1).$
Lemma 4.1.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary weakly $V$-geometrically ergodic transition operator, with $V,D,b$ defined as in Condition 3, and let $p\geq 1$ such that
$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty.$
Then,
$\mathcal{J}_{p}^{(s)}\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right).$
Proof.
Directly from the assumptions,
$\mathcal{J}_{p}^{(s)}$ $\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\pi(x)\left[\pi(x^{\prime})+\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]\right)^{1-1/p}$
$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\sum_{x,x^{\prime}\in\mathcal{X}}\left[\pi(x)\frac{1}{V(x^{\prime})}\left(D(x)+b(s-1)\right)\right]^{1-1/p}$
$\leq\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right),$
where the two inequalities stem from the sub-additivity of the function $x\mapsto x^{1-1/p}$. ∎
Corollary 4.1 (Under $V$-geometric ergodicity).
Let $P$ a $\pi$-stationary $V$-geometrically ergodic transition operator, with $V,M,\theta$ defined as in Condition 1. Then Condition 3 is satisfied with
$D=M\theta V\text{ and }b=0.$
If the following summability conditions are met for some $p\geq 1$,
$\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\pi\circ V\right\|_{1-1/p}^{1-1/p}<\infty,$
we obtain a control of $\mathcal{J}_{p}^{(s)}$ uniformly over $s$:
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\\ \leq\|\pi\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left\|\pi\circ D\right\|^{1-1/p}_{1-1/p}.$
Proof.
We first check Condition 3. For any $x,x^{\prime}\in\mathcal{X}$,
$V(x)(P^{s}(x,x^{\prime})-\pi(x^{\prime}))$ $\leq\sup_{x^{\prime\prime}\in\mathcal{X}}V(x^{\prime\prime})|P^{s}(x,x^{\prime\prime})-\pi(x^{\prime\prime})|$
$=\left\|e_{x}P^{s}-\pi\right\|_{V}\leq MV(x)\theta^{s}\leq MV(x)\theta,$
where the second inequality stems from the definition of $V$-geometric ergodicity. Thus, Condition 3 is satisfied with $D=M\theta V$ and $b=0$. An application of Lemma 4.1 gives the the second claim. ∎
Remark 4.1.
Under the assumption $D\propto V$, we can clarify the summability assumptions in Corollary 4.1.
  • Let us first assume that $\pi(x)\propto x^{-\alpha}$ for $\alpha\in(1,\infty)$. Then, for $V(x)\propto D(x)\propto x^{\beta}$, with $\beta\in\mathbb{R}_{+}$, $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\beta>p/(p-1)$, and $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ holds only when $\alpha-\beta>p/(p-1)$, thus we must have $p/(p-1)<\beta<\alpha-p/(p-1)$, which will only be possible for $\alpha>2p/(p-1)$.
  • Let us now assume now that $\pi(x)\propto a^{-x}$ for $a\in(0,1)$. Then, for $V(x)\propto D(x)\propto x^{\beta}$, $\left\|\pi\circ D\right\|_{1-1/p}^{1-1/p}<\infty$ always holds but we still require $\beta>p/(p-1)$ in order for $\left\|1/V\right\|_{1-1/p}^{1-1/p}<\infty$ to hold.
Lemma 4.2 (Under the existence of a dominating distribution).
Assume Condition 2 is satisfied. If moreover, for $p\geq 1$,
$\left\|\pi\right\|_{1-1/p}^{1-1/p}<\infty,\qquad\left\|\nu\right\|_{1-1/p}^{1-1/p}<\infty,$
then,
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq C^{1-1/p}\left\|\nu\right\|_{1-1/p}^{1-1/p}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
Proof.
It directly follows from the assumptions that for any $x,x^{\prime}\in\mathcal{X}$, and for any $s\in\mathbb{N}$,
$\begin{split}Q^{(s)}(x,x^{\prime})&=\pi(x)\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\\ &\leq C\pi(x)\nu(x^{\prime})\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s-1}(x,x^{\prime\prime})=C\pi(x)\nu(x^{\prime}).\end{split}$
As a result,
$\sum_{x,x^{\prime}\in\mathcal{X}}Q^{(s)}(x,x^{\prime})^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}\nu(x^{\prime})^{1-1/p}.$
In practice, Condition 3 can be readily verified through the following lemma.
Lemma 4.3.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Assume there exists functions $V\colon\mathcal{X}\to[1,\infty)$, $D\colon\mathcal{X}\to[1,\infty)$ and a constant $b\geq 0$ such that
$\sup_{x^{\prime}\in\mathcal{X}}V(x^{\prime})(P(x,x^{\prime})-\pi(x^{\prime}))\leq D(x)$
and, for all $x\in\mathcal{X}$, $PD(x)\leq D(x)+b$. Then, Condition 3 holds.
In Section 6, we use Lemma 4.3 to check Condition 3 on a whole family of transition operators $(P_{T})_{T\in\mathcal{T}}$. Note that Lemma 4.3 also shows that Condition 2 immediately implies Condition 3 with $D=C$, $V=1$ and $b=0$, which completes Figure 1.
Proof.
By recursion, for any $t\in\mathbb{N}$,
$\forall x\in\mathcal{X},\,P^{t}D(x)\leq D(x)+bt.$ (9)
Indeed, (9) holds for $t=1$ by assumption. Assume it holds for $t-1$, then
$P^{t}D(x)$ $=P[P^{t-1}D](x)$
$=\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[P^{t-1}D](x^{\prime})$
$\leq\sum_{x^{\prime}\in\mathcal{X}}P(x,x^{\prime})[D(x^{\prime})+b(t-1)]$
$=PD(x)+b(t-1)$
$\leq D(x)+bt.$
Thus, for any $x^{\prime\prime}\in\mathcal{X}$,
$V(x^{\prime\prime})[P^{t}(x,x^{\prime\prime})-\pi(x^{\prime\prime})]$ $=\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})[P(x^{\prime},x^{\prime\prime})-\pi(x^{\prime\prime})]V(x^{\prime\prime})$
$\leq\sum_{x^{\prime}\in\mathcal{X}}P^{t-1}(x,x^{\prime})D(x^{\prime})$
$=P^{t-1}D(x)$
$\leq D(x)+b(t-1),$
where the last inequality follows from (9). ∎
4.2   Uniformly bounded conditional entropy
When for each state $x\in\mathcal{X}$, the conditional distribution $e_{x}P$ has bounded entropy, it is possible to guarantee finiteness of $\mathcal{J}_{p,\xi}$.
Lemma 4.4.
Let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. Suppose that for $p\geq 1$,
$\mathcal{I}_{p}\doteq\sup_{x\in\mathcal{X}}\left\{\left\|e_{x}P\right\|^{1-1/p}_{1-1/p}\right\}<\infty.$
Then, for any $s\in\mathbb{N}$, it holds that
$\mathcal{J}_{p}^{(s)}\leq\mathcal{I}_{p}^{s}\left\|\pi\right\|_{1-1/p}^{1-1/p}.$
Proof.
Suppose that the claim holds for $s\geq 1$. Then,
$\begin{split}\mathcal{J}_{p}^{(s+1)}&=\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\left(\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})P(x^{\prime\prime},x^{\prime})\right)^{1-1/p}\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime\prime}\in\mathcal{X}}P^{s}(x,x^{\prime\prime})^{1-1/p}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &=\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P(x^{\prime\prime},x^{\prime})^{1-1/p}\\ &\leq\sum_{x,x^{\prime\prime}\in\mathcal{X}}\pi(x)^{1-1/p}P^{s}(x,x^{\prime\prime})^{1-1/p}\sup_{x\in\mathcal{X}}\left\|e_{x}P\right\|_{1-1/p}^{1-1/p}\\ &\leq\mathcal{J}_{p}^{(s)}\mathcal{I}_{p}.\end{split}$
The case $s=1$ is easy to verify, thus the lemma holds by induction. ∎
While a direct consequence of Lemma 4.4 is that
$\mathcal{J}_{p,\xi}\leq\mathcal{I}_{p}^{t^{\sharp}_{\mathsf{mix}}(\xi)}\left\|\pi\right\|_{1-1/p}^{1-1/p},$
guaranteeing finiteness of $\mathcal{J}_{p,\xi}$ for any $\xi\in(0,1)$, the above approach is only practical when bounding $\mathcal{J}_{p}^{(s)}$ for restrictively small values of $s$.
4.3   Infinite graphs with controlled growth
Natural structural assumptions on the underlying connection graph of a Markov chain also enable us to control the entropic term. Recall that an ergodic transition operator $P\in\mathcal{W}(\mathcal{X})$ induces a fully connected digraph $(\mathcal{X},\mathcal{E})$, where
$\mathcal{E}\doteq\left\{(x,x^{\prime})\in\mathcal{X}^{2}\colon P(x,x^{\prime})>0\right\}$
is the set of directed edges. The growth of a graph (Trofimov, 1985; Campbell et al., 2022) is defined as a function $g\colon\mathbb{N}\to\mathbb{N}\cup\left\{\infty\right\}$, where for $s\in\mathbb{N}$,
$g(s)\doteq\sup_{\mathcal{S}\subset\mathcal{X},\rho\left((\mathcal{X},\mathcal{E})|_{\mathcal{S}}\right)\leq s}\left|\mathcal{S}\right|,$
where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
Condition 4 (Connection graph with polynomial growth).
A graph is said to have at most polynomial growth when there exists $G\in\mathbb{R}_{+}$ and $q\geq 1$ such that for any $s\in\mathbb{N}$, it holds that $g(s)\leq Gs^{q}$. Then $q$ and $G$ are respectively called the degree and factor of growth.
Lemma 4.5.
Let $p,q\in[1,\infty),\xi\in(0,1)$ and let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. If the connection graph of $P$ has polynomial growth of degree $q$ with factor $G$, it holds that
$\mathcal{J}_{p,\xi}\leq G^{1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q/p}\left\|\pi\right\|^{1-1/p}_{1-1/p}.$
Proof.
For $x\in\mathcal{X}$, let us denote $\mathcal{S}_{x}^{(s)}\subset\mathcal{X}$ the set of states accessible from $x$ in $s$ steps. We have that
$\begin{split}\mathcal{J}_{p}^{(s)}&=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P^{s}(x,x^{\prime})^{1-1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{S}_{x}^{(s)}}\left(\frac{1}{\mathcal{S}_{x}^{(s)}}\right)^{1-1/p}\\ &=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}{\mathcal{S}_{x}^{(s)}}^{1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}(s^{q}G)^{1/p},\\ \end{split}$
where the first inequality stems from the uniform distribution maximizing the $\ell_{1-1/p}$ quasi-norm, and the second inequality follows from the polynomial growth assumption. ∎
Observe that
$\lim_{p\to\infty}\left\|\pi\right\|_{1-1/p}^{1-1/p}=\lim_{p\to\infty}\left[t^{\sharp}_{\mathsf{mix}}(\xi)\right]^{q/p}=1,$
that is, for large values of $p$, the upper bound of Lemma 4.5 still approaches unity.
Example 4.1 (Linear growth and stationary distribution with polynomially decaying tail).
We let $\mathcal{X}=\mathbb{N}$, and suppose that there exists $C\in\mathbb{R}_{+}$, and $\eta\in(1/8,1/4)$ such that $\pi(x)\leq\frac{C}{x^{1+\eta}}$ for any $x\in\mathcal{X}$. Taking $p=2(1+1/\eta)$, it holds that
$\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{(1+\eta)(1-1/p)}}=C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{1+\eta/2}}=C^{2\frac{\eta+2}{\eta+1}}\zeta(1+\eta/2),$
where $\zeta$ is the Riemann Zeta function. When the connection graph of $P$ enjoys linear growth, that is $q=1$, an application of Theorem 3.4, yields that for $\delta\in(0,1/2)$, a trajectory length of
$n\geq\widetilde{\mathcal{O}}\left(\frac{t_{\mathsf{mix}}^{11/10}}{\delta^{2}}\right),$
is sufficient to obtain $\widehat{t_{\mathsf{mix}}}^{\sharp}(2\delta/5)\geq t^{\sharp}_{\mathsf{mix}}(\delta/2)$ with probability at least $1-\delta$.
4.4   General finite space
When the state space $\mathcal{X}$ is finite, the graph-related methods discussed in the previous section remain applicable. Alternatively, a more direct approach yields the following bound,
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\frac{1}{\left|\mathcal{X}\right|^{2}}\right)^{1-1/p}=\left|\mathcal{X}\right|^{2/p},$ (10)
agnostic of the mixing properties of $P$. Note that (10) matches the heuristic lower bound $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ in the worst-case where $P$ is bistochastic.
Corollary 4.2 (of Theorem 3.4 for $\left|\mathcal{X}\right|<\infty$).
Let $\mathcal{X}$ be a finite space. Let $\varepsilon\in(0,1/4)$, and $\delta,\xi\in(0,1)$. Let $n\in\mathbb{N}$, and let $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\colon\mathcal{X}^{n}\to\mathbb{N}$ be the estimator for $t^{\sharp}_{\mathsf{mix}}(\xi)$ defined in (6). For any ergodic stationary Markov chain $X_{1},X_{2},\dots,X_{n}$ with transition operator $P\in\mathcal{W}(\mathcal{X})$. When
$n\geq 1+\frac{2^{8}et_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+2\log_{2}\left|\mathcal{X}\right|\right),9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
with probability at least $1-\delta$, it holds that
$\begin{split}\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)\in\left[t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon)),t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))\right].\end{split}$
Proof.
We plug (10) in Theorem 3.4 and optimize for $p$. The function $p\mapsto p\left|\mathcal{X}\right|^{2/p}$ is minimized for $p=2\log\left|\mathcal{X}\right|$, whence the corollary. ∎
4.5   The two-point space
In this section, we analyze the case where $\left|\mathcal{X}\right|=2$, illustrating the possibly stark difference between the worst-case mixing time and the average-mixing time. For $p,q\in(0,1)$, we let
$P_{p,q}=\begin{pmatrix}1-p&p\\ q&1-q\end{pmatrix}.$
Quantities such as $d$ or $d^{\sharp}$ subscripted with $p,q$ will be understood as pertaining to $P_{p,q}$. The matrix $P_{p,q}$ is irreducible, and its stationary distribution is readily given by
$\pi_{p,q}^{\intercal}=\frac{1}{p+q}\begin{pmatrix}q\\ p\end{pmatrix}.$
It follows from a direct computation (see e.g. the proof of Wolfer (2023+, Lemma 8.1)) that
$\begin{split}P_{p,q}^{t}&=1^{\intercal}\pi_{p,q}+\frac{(1-p-q)^{t}}{p+q}\begin{pmatrix}p&-p\\ -q&q\end{pmatrix},\\ \end{split}$
and as a result,
$\begin{split}d_{p,q}^{\sharp}(t)&=2\frac{pq}{(p+q)^{2}}\left|1-p-q\right|^{t},\qquad d_{p,q}(t)=\frac{p\lor q}{p+q}\left|1-p-q\right|^{t}.\end{split}$
When $p=q$, the stationary distribution is uniform, $d_{p,q}^{\sharp}(t)=d_{p,q}(t)$, and
$t_{\mathsf{mix}}(\xi)=t_{\mathsf{mix}}^{\sharp}(\xi).$
However, for $\eta\doteq p/q<1$, the ratio of the distances is a function of the parameter $\eta$,
$\frac{d_{p,q}(t)}{d^{\sharp}_{p,q}(t)}=\frac{(1+\eta)^{2}}{2\eta}\doteq r_{\eta}(t),$
whence the following lemma holds.
Lemma 4.6.
There exists a family of Markov chains
$\left\{P_{p,q}\colon(p,q)\in(0,1)^{2}\right\}\in\mathcal{W}^{\star}(\left\{0,1\right\})$
such that for any $\xi\in(0,1)$ and any $M>0$, there exist $p,q\in(0,1)$ such that
$t_{\mathsf{mix}}(\xi)>Mt_{\mathsf{mix}}^{\sharp}(\xi).$
5   Proofs
In this section, we compile technical proofs that were either outlined briefly or omitted in the manuscript. The subsequent technical tool will prove to be convenient.
Lemma 5.1.
Let $A,B\in\mathbb{R}_{+}$, with $A,B\geq e$.
$B\geq 2A\log A\implies B\geq A\log B.$
Proof.
The function $x\mapsto x/\log x$ is increasing on $[e,\infty)$ and in this range, it holds that $x/\log x\geq e>2$. Hence, for $B\geq 2A\log A$, it holds that
$\frac{B}{\log B}\geq\frac{2A\log A}{\log A+\log(2\log A)}\geq\frac{2A\log A}{\log A+\log A}=A,$
whence the lemma holds. ∎
5.1   Proof of Lemma 2.3
The proof is standard, and based on the blocking technique credited to Bernstein and further developed in the series of papers (Yu, 1993, 1994; Eberlein, 1984; Volkonskii and Rozanov, 1959). For simplicity, we assume that $n=2Bs$, where $B,s\in\mathbb{N}$. For $b\in[B]$, we write
$\begin{split}X^{[2b]}&=X^{[2b]}_{1},\dots,X^{[2b]}_{s}=X_{(2b-1)s+1},\dots,X_{2bs},\\ X^{[2b-1]}&=X^{[2b-1]}_{1},\dots,X^{[2b-1]}_{s}=X_{(2b-2)s+1},\dots,X_{(2b-1)s},\\ \end{split}$
and consider the partitioned process
$X=X^{[1]},X^{[2]},\dots,X^{[2B]}.$
In other words, we decompose the observed trajectory into $B$ even and $B$ odd blocks of size $s$, as illustrated in Figure 2.
{tikzpicture}
Figure 2: Decomposing the process into blocks.
Invoking Hoeffding’s inequality (Hoeffding, 1963) and Yu (1994, Corollary 2.7), it holds that
$\begin{split}\mathbb{P}\left(\frac{1}{n}\sum_{t=1}^{n}f(X_{t})>\varepsilon\right)&\leq\sum_{\sigma\in\left\{0,1\right\}}\mathbb{P}\left(\sum_{b=1}^{B}\sum_{t=1}^{s}f\left(X_{t}^{[2b-\sigma]}\right)>n\varepsilon/2\right)\\ &\leq 2\exp\left(-\frac{n\varepsilon^{2}}{4s}\right)+2\left(B-1\right)\beta(s).\end{split}$ (11)
We first prove the claim in the sub-exponential $\beta$-mixing setting, that is for $\beta(s)\leq\beta_{0}e^{-\beta_{1}s^{b}}$ with parameters $b,\beta_{0},\beta_{1}$ in their specified range. For any $\xi\in(0,1)$, it holds that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\left\lceil\left(\frac{1}{\beta_{1}}\log\frac{\beta_{0}}{\xi}\right)^{1/b}\right\rceil.$
Observing that
$2(B-1)\beta(s)\leq\frac{n\beta_{0}}{s}\exp(-\beta_{1}s^{b}),$
and denoting $W_{0}$ the principal branch of the product logarithm Lambert function (Corless et al., 1996), for
$s=\left\lceil\left(\frac{1}{b\beta_{1}}W_{0}\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right)\right)^{1/b}\right\rceil,$
it holds that $22(B-1)\beta(s)\leq\delta/$ 2. As a consequence, the deviation probability at (11) is bounded by $\delta$ when the trajectory length $n$ satisfies the inequality,
$n\geq\frac{4}{\varepsilon^{2}}\left\lceil\left(\frac{1}{b\beta_{1}}W_{0}\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right)\right)^{1/b}\right\rceil\log\left(\frac{4}{\delta}\right).$
We now give an approximate albeit more tractable sufficient condition on $n$. Since $b\beta_{1}\leq 1$, for the above argument of $W_{0}$ to be no smaller than $e$, it is enough to have,
$n\geq\frac{\delta}{2\beta_{0}}\left(\frac{e}{b\beta_{1}}\right)^{1/b}.$
Since $\beta_{0}\geq 1,b\leq 1$ and $x\leq\frac{1}{e}\implies x\leq\log\frac{1}{x}$,
$\frac{\delta}{2\beta_{0}}\left(\frac{e}{b\beta_{1}}\right)^{1/b}\leq\frac{e}{2}\log\left(\frac{e\beta_{0}}{\delta}\right)\left(\frac{e}{b\beta_{1}}\right)^{1/b}\leq\frac{e}{2}\left(\frac{e}{b\beta_{1}}\log\left(\frac{e\beta_{0}}{\delta}\right)\right)^{1/b}\leq\frac{e}{2}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}(\delta/e).$
Therefore,
$n\geq\frac{e}{2}\left(\frac{e}{b}\right)^{1/b}t^{\sharp}_{\mathsf{mix}}(\delta/e),$
is sufficient. Further relying on the upper bounds
$x\geq e\implies W_{0}(x)\leq\log(x)\;\text{and}\;x\geq 1\implies\lceil x\rceil\leq 2x,$
it suffices for $n$ to satisfy
$n^{b}\geq\left(\frac{8}{\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)\right)^{b}\frac{1}{b\beta_{1}}\log\left(b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b}\right).$
Setting
$A=\left[\frac{16\beta_{0}}{\delta\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)\right]^{b},\qquad B=b\beta_{1}\left[\frac{2n\beta_{0}}{\delta}\right]^{b},$
and invoking Lemma 5.1, it is enough that
$n\geq\frac{8}{\varepsilon^{2}}\log\left(\frac{4}{\delta}\right)2^{1/b}t^{\sharp}_{\mathsf{mix}}\left(\xi(\varepsilon,\delta)\right),$
with $\xi(\varepsilon,\delta)=\frac{\delta\varepsilon^{2}}{16\log\left(\frac{4}{\delta}\right)}$. In order to streamline this upper bound, observe that
$\frac{1}{\beta_{1}}\log\left(\frac{\beta_{0}}{\xi(\varepsilon,\delta)}\right)\leq t^{\sharp}_{\mathsf{mix}}(\delta/e)+t^{\sharp}_{\mathsf{mix}}(\delta/e)\log\left(\frac{16\log(4/\delta)}{e\varepsilon^{2}}\right),$
and recall that for $b\in(0,1]$
$x,y\in\mathbb{R}_{+}\implies(x+y)^{1/b}\leq\frac{1}{b}\left(x^{1/b}+y^{1/b}\right).$
Finally, let us now move on to the geometric case, that is,
$\beta(s)\leq\frac{\beta_{1}}{s^{b}},$
for some $\beta_{1},b>0$. For any $\xi\in(0,1)$, it holds that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq\left\lceil\left(\frac{\beta_{1}}{\xi}\right)^{1/b}\right\rceil.$
Having $\beta_{1}\geq 1$ ensures that $\frac{2n\beta_{1}}{\delta}\geq 1$. Setting
$s=\left\lceil\left(\frac{2n\beta_{1}}{\delta}\right)^{\frac{1}{b+1}}\right\rceil,$
yields
$\frac{n}{s}\beta(s)\leq\frac{n\beta_{1}}{s^{b+1}}\leq n\beta_{1}\left(\frac{\delta}{2n\beta_{1}}\right)=\frac{\delta}{2}.$
Finally, setting
$n\geq\left(\frac{8}{\varepsilon^{2}}\log\frac{4}{\delta}\right)^{\frac{b+1}{b}}\left(\frac{2\beta_{1}}{\delta}\right)^{\frac{1}{b}}$
leads to
$2\exp\left(-\frac{n\varepsilon^{2}}{4s}\right)\leq 2\exp\left(-\frac{n\varepsilon^{2}}{8\left(\frac{2n\beta_{1}}{\delta}\right)^{\frac{1}{b+1}}}\right)\leq\frac{\delta}{2}.$
For both the sub-exponential and the geometric case, an adaption of the proof handles the general case where $n$ is not divisible by the block size $s$, modulo paying a universal multiplicative constant. ∎
5.2   Proof of Lemma 2.1
The proof closely follows the techniques in Levin et al. (2009, Lemma 12.2, Theorem 12.4), albeit in a countably infinite alphabet setting. The linear operator defined by $P$ is self-adjoint in $\ell_{2}(\pi)$, hence normal. The assumption on compactness of $P$ thus enables diagonalization methods. Suppose that $P$ is reversible. We can write,
$\begin{split}d^{\sharp}(t)&\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|\delta_{x}P^{t}-\pi\right\|_{\mathsf{TV}}\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\left|P^{t}(x,x^{\prime})-\pi(x^{\prime})\right|\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\\ &\stackrel{{\scriptstyle(\spadesuit)}}{{\leq}}\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\frac{(1-\gamma_{\star})^{t}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\sum_{x,x^{\prime}\in\mathcal{X}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\left\|\pi\right\|_{1/2}.\\ \end{split}$
For the inequality ($\spadesuit$), we used compactness of $P$ and the fact Levin et al. (2009, Lemma 12.2) that for the $\langle\cdot,\cdot\rangle_{\pi}$-orthonormal basis of real-valued eigenfunctions $(f_{i})_{x\in\mathcal{X}}$,
$\begin{split}P^{t}(x,x^{\prime})&=\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x^{\prime})\pi(x^{\prime})\lambda^{t}_{i}f_{i}(x)=\pi(x^{\prime})+\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}(x^{\prime})\pi(x^{\prime})\lambda^{t}_{i}f_{i}(x)\\ \delta_{x}&=\sum_{i=1}^{\left|\mathcal{X}\right|}\langle\delta_{x},f_{i}\rangle_{\pi}f_{i}=\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)\pi(x)f_{i}\end{split}$
$\pi(x)=\langle\delta_{x},\delta_{x}\rangle_{\pi}=\Biggl{\langle}\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)\pi(x)f_{i},\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)\pi(x)f_{i}\Biggr{\rangle}_{\pi}=\pi(x)^{2}\sum_{i=1}^{\left|\mathcal{X}\right|}f_{i}(x)^{2},$
thus as in Levin et al. (2009, Theorem 12.4),
$\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\leq\left|\lambda_{\star}\right|^{t}\sqrt{\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x)\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x^{\prime})}.$
Solving for a proximity parameter $\xi\in(0,1)$, we obtain that
$t^{\sharp}_{\mathsf{mix}}(\xi)\leq t_{\mathsf{rel}}\log\frac{\left\|\pi\right\|_{1/2}}{2\xi}.$
5.3   Proof of Theorem 3.1
For $s\in\mathbb{N}$,
$\begin{split}2\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|=&\mathbb{E}_{\pi}\left|\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}-\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor}\right|-\sum_{x\in\mathcal{X}}\pi(x)\left\|e_{x}P^{s}-\pi\right\|_{1}\right|\\ \leq&\mathbb{E}_{\pi}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N_{xx^{\prime}}^{(s)}}{\lfloor(n-1)/s\rfloor}-\pi(x)P^{s}(x,x^{\prime})\right|\\ &+\mathbb{E}_{\pi}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|.\end{split}$
We further decompose the argument in the second expectation as
$\begin{split}&\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}\right|+\left|\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}-\pi(x)\pi(x^{\prime})\right|\\ &=2\sum_{x\in\mathcal{X}}\left|\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}-\pi(x)\right|.\\ \end{split}$
Noting that by stationarity,
$\mathbb{E}_{\pi}\left[\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}\right]=\pi(x),\qquad\mathbb{E}_{\pi}\left[\frac{N_{xx^{\prime}}^{(s)}}{\lfloor(n-1)/s\rfloor}\right]=\pi(x)P^{s}(x,x^{\prime}),$
it follows from Jensen’s inequality that
$\begin{split}2\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq\frac{1}{\lfloor(n-1)/s\rfloor}\left(\sum_{x,x^{\prime}\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]}+2\sum_{x\in\mathcal{X}}\sqrt{\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]}\right).\end{split}$ (12)
We proceed to bound the two variance terms. We first treat $\operatorname{Var}_{\pi}\left[N_{x}\right]$ for a fixed $x\in\mathcal{X}$. This will in turn yield bounds for general $\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]$ and $\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]$. The sequence $(\delta\left[X_{t}=x\right])_{t\in\mathbb{N}}$ being stationary at second order, we expand the variance of the sum (see e.g. Rio (1999, (1.6))) as,
$\operatorname{Var}_{\pi}\left[N_{x}\right]=(n-1)\operatorname{Var}_{\pi}\left[\delta\left[X_{1}=x\right]\right]+2\sum_{t=1}^{n-2}(n-1-t)\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right]).$ (13)
From Rio (1993, Theorem 1) [^7] This generalization of Ibragimov’s covariance inequality (Ibragimov, 1962) also appears in Davydov (1968) but with a weaker constant. (see also Rio (1999, (1.12b)) and Doukhan (2012, Theorem 3)), for a Hölder triplet $p,r,q\geq 1,\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$ and any $t\in\mathbb{N}\cup\left\{0\right\}$,
$\begin{split}\operatorname{Cov}_{\pi}(\delta\left[X_{1}=x\right],\delta\left[X_{1+t}=x\right])&\leq 2(2\alpha(t))^{1/p}\left\|\delta\left[X_{1}=x\right]\right\|_{\pi,r}\left\|\delta\left[X_{t}=x\right]\right\|_{\pi,q}\\ &\leq 2\beta(t)^{1/p}\pi(x)^{1-1/p},\end{split}$
where $\alpha$ is Rosenblatt’s strong mixing coefficient [^8] We note that $\alpha$ is defined as twice Rosenblatt’s coefficient in Rio (1999). (Rosenblatt, 1956), we wrote $\left\|X\right\|_{\pi,r}=\left(\mathbb{E}_{\pi}\left|X\right|^{r}\right)^{1/r}$ for a real random variable $X$, and for the second inequality, we relied on the inequality $2\alpha(t)\leq\beta(t)$ (see e.g. (Bradley, 2005, (1.11))). Plugging in (13), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 2(n-1)\beta(0)^{1/p}\pi(x)^{1-1/p}+4\sum_{t=1}^{n-2}(n-1-t)\beta(t)^{1/p}\pi(x)^{1-1/p}\\ &\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right),\end{split}$
and for a general $s<n$, noting that the mixing coefficient $\beta^{(s)}(t)$ pertaining to the $s$-skipped chain satisfies $\beta^{(s)}(t)=\beta(st)$, we obtain
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p}\right)\\ &\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\mathcal{B}^{(s)}_{p}.\\ \end{split}$
We verify (see e.g. Wolfer and Kontorovich (2021, Lemma 6.1)) that the process consisting of a sliding window over pairs,
$\{E_{t}\}_{t\in\mathbb{N}}\doteq\left\{\left(X_{t},X_{t+1}\right)\right\}_{t\in\mathbb{N}}$
forms a Markov chain over the state space =$\mathcal{E}=\left\{x,x^{\prime}\in\mathcal{X}\colon P(x,x^{\prime})>0\right\}$, called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2022) of the original chain, with transition operator
$\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),$
and with stationary distribution $\overline{\pi}((x,x^{\prime}))=\pi(x)P(x,x^{\prime})$. Writing $\overline{\beta}(s)$ for the $\beta$-mixing coefficient of $\overline{P}$, we compute
$\begin{split}\overline{\beta}(0)&=1-\left\|\overline{\pi}\right\|_{2}^{2}\\ \overline{\beta}(1)&=\beta(0)=1-\left\|\pi\right\|_{2}^{2},\end{split}$
and it follows from the Markov property that for $t\in\mathbb{N},t<n-1$,
$\overline{P}^{t+1}((x_{1},x_{2}),(x_{1}^{\prime},x_{2}^{\prime}))=P(x_{1}^{\prime},x_{2}^{\prime})P^{t}(x_{2},x_{1}^{\prime}),$
and thus, for $t\geq 1$,
$\begin{split}\overline{\beta}(t+1)&=\frac{1}{2}\sum_{e=(x_{1},x_{2})\in\mathcal{E}}\overline{\pi}(e)\sum_{e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime})\in\mathcal{E}}\left|\overline{P}^{t+1}(e,e^{\prime})-\overline{\pi}(e^{\prime})\right|\\ &=\frac{1}{2}\sum_{(x_{1},x_{2})\in\mathcal{E}}\pi(x_{1})P(x_{1},x_{2})\sum_{(x^{\prime}_{1},x^{\prime}_{2})\in\mathcal{E}}\left|P^{t}(x_{2},x_{1}^{\prime})-\pi(x_{1}^{\prime})\right|P(x_{1}^{\prime},x_{2}^{\prime})\\ &=\frac{1}{2}\sum_{x_{2}\in\mathcal{X}}\pi(x_{2})\sum_{x^{\prime}_{1}\in\mathcal{X}}\left|P^{t}(x_{2},x_{1}^{\prime})-\pi(x_{1}^{\prime})\right|=\beta(t).\end{split}$
More generally, for $s\in\mathbb{N},t\in\mathbb{N}\cup\left\{0\right\}$, $1\leq s,\leq n-1$, and $0\leq t<\lfloor(n-1)/s\rfloor$,
$\overline{\beta}^{(s)}(t+1)=\beta^{(s)}(t)=\beta(st).$
As a consequence
$\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor Q^{(s)}(x,x^{\prime})^{1-1/p}\left(\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}\right),$
where we recall that $Q^{(s)}(x,x^{\prime})=\pi(x)^{s}P(x,x^{\prime})$. We further bound
$\begin{split}\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}&=\frac{1}{2}\left(1-\left\|Q^{(s)}\right\|_{2}^{2}\right)+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(s(t-1))^{1/p}\\ &\leq\frac{1}{2}\left(1-\left\|\pi\right\|_{2}^{2}\right)+\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st))^{1/p}\\ &\leq\frac{3}{2}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st)^{1/p}\leq\frac{3}{2}\mathcal{B}_{p}^{(s)},\end{split}$
and plugging in (12),
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\frac{3}{2}\mathcal{B}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}\mathcal{B}_{p}^{(s)}}\right),\end{split}$
which yields the first inequality in the theorem. The second inequality can be obtained from observing that $\left\|Q^{(s)}\right\|_{1-1/p}\geq\left\|\pi\right\|_{1-1/p}$. ∎
5.4   Proof of Corollary 3.1
Let us first analyze the exponential case, by assuming that $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s)$.
$\begin{split}\mathcal{B}_{p}^{(s)}=\beta_{0}^{1/p}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\exp\left(-\beta_{1}st/p\right)&\leq\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}.\end{split}$
We proceed to address the sub-exponential case. Let us assume that $\beta(s)\leq\beta_{0}\exp(-\beta_{1}s^{b})$ for $\beta_{0}>0,\beta_{1}>0$ and $b\in(0,1)$. Then,
[…]
Finally, in the polynomial case, where for $s\geq 1$, $\beta(s)\leq\beta_{1}/s^{b}$,
$\begin{split}\mathcal{B}_{p}^{(s)}&\leq\beta_{0}^{1/p}+\beta(s)^{1/p}\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\frac{1}{t^{b/p}}\leq\beta_{0}^{1/p}+\zeta(b/p)\beta(s)^{1/p}.\\ \end{split}$
5.5   Proof of Theorem 3.2
Recall that for a uniformly ergodic Markov chain,
$\beta(s)\leq d(s)\leq d(\lfloor s/t_{\mathsf{mix}}\rfloor t_{\mathsf{mix}})\leq\exp(-\lfloor s/t_{\mathsf{mix}}\rfloor\log 2)\leq 2\exp\left(-\frac{\log 2}{t_{\mathsf{mix}}}s\right),$
where the third inequality is adapted from Levin et al. (2009, (4.33)), i.e. $\beta$ is decays exponentially with $\beta_{0}=2$ and $\beta_{1}=\log(2)/t_{\mathsf{mix}}$. Thus, from the fact that $x>0\implies 1/(1-\exp(-x))\leq 1+1/x,$ and Corollary 3.1, it holds that
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq 4\sqrt{\frac{2^{1/p}}{\lfloor(n-1)/s\rfloor}\left(1+\frac{pt_{\mathsf{mix}}}{s\log 2}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}},\\ \end{split}$
and $(i)$ follows from $n>s+1$ and $x>1\implies\lfloor x\rfloor\geq x/2$ and $2^{1/p}\leq 2$. From $(i)$, when
$n\geq 1+\frac{2^{8}}{\varepsilon^{2}}\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p},$
it holds that, $\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\leq\varepsilon/2$ and in this case, the probability of the above deviation can be upper bounded by fluctuations around the mean as follows,
$\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq\mathbb{P}_{\pi}\left(\left|\left|\widehat{\beta}(s)-\beta(s)\right|-\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|\right|>\varepsilon/2\right).$
Let
$x_{[s]}^{n}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+\lfloor(n-1)/s\rfloor s}\in\mathcal{X}$
be some $s$-skipped trajectory, let $t\in\left\{0,\dots,\lfloor(n-1)/s\rfloor\right\}$, $x^{\prime}_{1+ts}\in\mathcal{X}$, and denote
$x^{n,(t)}=x_{1},x_{1+s},x_{1+2s},\dots,x_{1+(t-1))s},x^{\prime}_{1+ts},x_{1+(t+1)s}\dots,x_{1+\lfloor(n-1)/s\rfloor s}$
the trajectory where the observation $x_{1+ts}$ at time $1+ts$ was replaced with observation $x^{\prime}_{1+ts}$. We show that the estimator satisfies the bounded differences condition. We decompose,
$\begin{split}&\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\\ &\leq\left|\widehat{\beta}(s)(X)-\widehat{\beta}(s)(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{xx^{\prime}}^{(s)}(X)-N_{xx^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|.\\ \end{split}$
We further upper bound the second term as follows,
$\begin{split}&\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X)-N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &+\frac{1}{\lfloor(n-1)/s\rfloor^{2}}\sum_{x,x^{\prime}\in\mathcal{X}}\left|N_{x}^{(s)}(X)N_{x^{\prime}}^{(s)}(X^{(t)})-N_{x}^{(s)}(X^{(t)})N_{x^{\prime}}^{(s)}(X^{(t)})\right|\\ &\leq\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x^{\prime}\in\mathcal{X}}\left|N_{x^{\prime}}^{(s)}(X)-N_{x^{\prime}}^{(s)}(X^{(t)})\right|+\frac{1}{\lfloor(n-1)/s\rfloor}\sum_{x\in\mathcal{X}}\left|N_{x}^{(s)}(X)-N_{x}^{(s)}(X^{(t)})\right|.\\ \end{split}$
We obtain
$\left|\left|\widehat{\beta}(s)(X)-\beta(s)\right|-\left|\widehat{\beta}(s)(X^{(t)})-\beta(s)\right|\right|\leq\frac{4}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}+\frac{2}{\lfloor(n-1)/s\rfloor}.$
From a version of McDiarmid’s inequality for uniformly ergodic Markov chains (Paulin, 2015, Corollary 2.10), and the fact that the mixing time of the $s$-skipped chain $t_{\mathsf{mix}}^{(s)}$ verifies $t_{\mathsf{mix}}^{(s)}\leq\lceil t_{\mathsf{mix}}/s\rceil$,
$\begin{split}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)&\leq 2\exp\left(-\frac{2(\varepsilon/2)^{2}}{9t_{\mathsf{mix}}^{(s)}\lfloor(n-1)/s\rfloor\left(\frac{8}{\lfloor(n-1)/s\rfloor}\right)^{2}}\right)\\ &\leq 2\exp\left(-\frac{(n-1)\varepsilon^{2}}{9\cdot 2^{8}(t_{\mathsf{mix}}+s)}\right),\end{split}$
which completes the proof of claim $(ii)$. ∎
5.6   Proof of Theorem 3.3
Suppose that
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},18\log\left(\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right)\right\}.$
Invoking Lemma 5.1 with
$a\doteq\frac{9\cdot 2^{10}}{\delta\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon),\qquad b\doteq\frac{2}{\delta_{n-1}},$
and where $\delta_{n-1}\doteq\frac{\delta}{n-1}$, it holds that
$n-1\geq\frac{2^{9}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}.$ (14)
Writing $S=S(n,\varepsilon,\delta,p,\mathcal{J}_{p})$, we thus have
$\begin{split}n-1&\geq\frac{n-1}{2}+\frac{2^{8}}{\varepsilon^{2}}\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\left[S+\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\right]\max\left\{\mathcal{J}_{p},9\log\frac{2}{\delta_{n-1}}\right\}\\ &\geq\frac{2^{8}}{\varepsilon^{2}}\max\left\{\left(S+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\mathcal{J}_{p},9(S+t_{\mathsf{mix}})\log\frac{2}{\delta_{n-1}}\right\}.\end{split}$
Furthermore, observing from (14) that
$S\geq\frac{p}{\log 2}t_{\mathsf{mix}}(\varepsilon)\geq t^{\sharp}_{\mathsf{mix}}(\varepsilon),$
it follows that $\beta(s)<\varepsilon$ for $s>S$. An application of the union bound and Theorem 3.2$(ii)$ yield that
$\mathbb{P}_{\pi}\left(\left\|\widehat{\beta}-\beta\right\|_{\infty}>\varepsilon\right)\leq\sum_{s=1}^{S}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|>\varepsilon\right)\leq S\delta_{n-1}\leq\delta.$
The claim then follows from $t_{\mathsf{mix}}(\varepsilon)\leq\lceil\log_{2}\varepsilon^{-1}\rceil t_{\mathsf{mix}}$ for any $\varepsilon\in(0,1/4)$ (Levin et al., 2009, (4.34)). ∎
5.7   Proof of Theorem 3.4
We first bound the probability that the estimator overshoots, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)>t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$. On this event, for any $s\leq t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, and in particular for $s=t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon))$, it must be that $\widehat{\beta}(s)>\xi$. From Theorem 3.2$(ii)$, for
$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4}{\delta}\right\},$
it holds that
$\begin{split}\mathbb{P}_{\pi}\left(\widehat{\beta}(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))>\xi\right)&\leq\mathbb{P}_{\pi}\left(\widehat{\beta}(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))-\beta(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))>\xi-\xi(1-\varepsilon)\right)\\ &\leq\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))-\beta(t_{\mathsf{mix}}^{\sharp}(\xi(1-\varepsilon)))\right|>\xi\varepsilon\right)\leq\frac{\delta}{2}.\end{split}$
As a second step, we bound the probability that $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)$ underestimates the true parameter, that is $\widehat{t_{\mathsf{mix}}}^{\sharp}(\xi)<t_{\mathsf{mix}}^{\sharp}(\xi(1+\varepsilon))$. From Theorem 3.2$(ii)$ and an application of the union bound, when for any $s\leq t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))$ the trajectory length satisfies
$n\geq 1+\frac{2^{8}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(s+\frac{p}{\log 2}t_{\mathsf{mix}}\right)\left\|Q^{(s)}\right\|^{1-1/p}_{1-1/p},9(s+t_{\mathsf{mix}})\log\frac{4t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))}{\delta}\right\},$
it holds that
$\begin{split}\mathbb{P}_{\pi}\left(\exists s<t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon)),\widehat{\beta}(s)\leq\xi\right)&\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\widehat{\beta}(s)\leq\xi\right)\\ &\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\beta(s)-\widehat{\beta}(s)\geq-\xi+\xi(1+\varepsilon)\right)\\ &\leq\sum_{s=1}^{t^{\sharp}_{\mathsf{mix}}(\xi(1+\varepsilon))-1}\mathbb{P}_{\pi}\left(\left|\widehat{\beta}(s)-\beta(s)\right|\geq\xi\varepsilon\right)\leq\frac{\delta}{2}.\end{split}$
The theorem follows from $t^{\sharp}_{\mathsf{mix}}(\xi(1-\varepsilon))\leq t_{\mathsf{mix}}(\xi(1-\varepsilon))\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil\leq t_{\mathsf{mix}}\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil$. ∎
6   Example
In this section, we identify $\mathcal{X}=\mathbb{N}$, and we study the properties of a family of transition operators $(P_{T})_{T\in\mathcal{T}}$.
Definition 6.1.
Let $q\in(0,1)$, a sequence $\nu=(\nu_{x})_{x=1,\dots}$ taking values in $[0,1]$, and a probability distribution $\mu=(\mu_{x})_{x=2,\dots}$. We define:
$P_{T}=\left(\begin{array}[]{c c c c c}1-q&q\mu_{2}&q\mu_{3}&q\mu_{4}&\dots\\ \nu_{2}&1-\nu_{2}&0&0&\dots\\ \nu_{3}&0&1-\nu_{3}&0&\dots\\ \nu_{4}&0&0&1-\nu_{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{array}\right).$
We define $\mathcal{T}$ as the set of triples $T=(q,\mu,\nu)$ such that $q>0$, $\nu_{x}>0$ and $\mu_{x}>0$ for all $x$ and $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<+\infty$.
A first observation is that when $q>0$ and $\nu_{x}>0$ and $\mu_{x}>0$ for all $x$, then the chain is aperiodic and irreducible. Thus it has an invariant positive measure (not necessarily finite). Solving $\pi P_{T}=\pi$, we find that the invariant measure satisfies, for $x\geq 2$,
$\pi(x)=\frac{q\mu_{x}}{\nu_{x}}\pi(1).$
So, if $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}=+\infty$, the chain is recurrent null. But if $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<\infty$ then it is recurrent positive, and the unique invariant probability measure, that will be denoted by $\pi_{T}$, is given by
$\pi_{T}(1)=\frac{1}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\text{ and }\pi_{T}(x)=\frac{q\mu_{x}}{\nu_{x}\left(1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}\right)}\text{ for }x>1.$ (15)
Thus, when $T=(q,\mu,\nu)\in\mathcal{T}$ the chain is recurrent positive. At first, we will study the ergodic properties of $P_{T}$ and evaluate its $\beta$-mixing coefficients. Secondly, we will apply the results of Section 4 to control $\mathcal{J}_{p}^{(s)}$ for $P_{T}$.
6.1   Ergodic and mixing properties of $P_{T}$
We remind that the Dobrushin coefficient of $P_{T}$ is defined by
$\kappa:=\sup_{x,y\in\mathbb{N}}\|P_{T}(x,\cdot)-P_{T}(y,\cdot)\|_{{\rm TV}}.$ (16)
An application of Theorem 18.2.4 in Douc et al. (2018) gives, for any $t\in\mathbb{N}$,
$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\leq\kappa^{t}.$ (17)
Thus, when $\kappa<1$, the chain is uniformly ergodic.
Note for example that if $q<1$ and $\nu_{x}\geq 1-q$, then for any $x$ we have $P_{T}(x,1)\geq 1-q$. Thus, for any $x$ and $y$, $\|P_{T}(x,\cdot)-P_{T}(y,\cdot)\|_{{\rm TV}}\leq q<1$, and thus $\kappa<q<1$. So the chain defined by $P_{T}$ is uniformly ergodic. In this case, $t_{\mathsf{mix}}(\xi)\leq\lceil\log(\xi)/\log q\rceil$ and the chain is exponentially $\beta$-mixing.
A first interesting observation is that when $\nu_{x}\rightarrow 0$, the chain cannot be uniformly ergodic.
Lemma 6.1.
Assume $T=(q,\mu,\nu)\in\mathcal{T}$. If $\nu_{x}\rightarrow 0$ for $x\rightarrow\infty$, we have
$d(t)=\sup_{x\geq 1}\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\geq\frac{1}{2}$
and thus $t_{{\rm mix}}=+\infty$.
Proof: we have, for a fixed $t$,
$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\sum_{y=1}^{\infty}|P_{T}^{t}(x,y)-\pi_{T}(y)|$
$\geq\frac{1}{2}|P_{T}^{t}(x,x)-\pi_{T}(x)|$
$\geq\frac{1}{2}|(1-\nu_{x})^{t}-\pi_{T}(x)|.$
Note that $\pi_{T}(x)\rightarrow 0$ for $x\rightarrow\infty$, while $(1-\nu_{x})^{t}\rightarrow 1$. $\square$
Of course, even though it is not uniformly ergodic, as the chain remains ergodic for $T\in\mathcal{T}$. Thus, for $t\rightarrow+\infty$ we have $\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\rightarrow 0$. Thus, by the dominated convergence theorem, in the stationary case, we have $\beta(t)=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\rightarrow 0.$ We can actually provide more accurate information on $\beta(t)$. This requires a careful analysis that will be simplified by additional technical assumptions on $T$.
Definition 6.2.
We define $\mathcal{S}\subset\mathcal{T}$ as the set of triples $T=(q,\mu,\nu)\in\mathcal{T}$ such that $q\in(0,1/4]$, $(\nu_{x})_{x=1,\dots}$ is nonincreasing and takes values in $(0,1/2]$ and $\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}<+\infty$.
Lemma 6.2.
Assume $T=(q,\mu,\nu)\in\mathcal{S}$. For any $K$ large enough to ensure $\sum_{x>K}\pi_{T}(x)\leq 1/2$,
$\beta(t)\leq 2(1-\nu_{K})^{t}+7\sum_{x>K}\pi_{T}(x)+2tq\sum_{x>K}\mu_{x}.$
We can simplify a little the above by noting that
$q\mu_{x}=\pi_{T}(x)\nu_{x}\left(1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}\right)\leq\frac{\pi_{T}(x)\nu_{x}}{\pi_{T}(1)}\leq\frac{\pi_{T}(x)}{2\pi_{T}(1)}$
and thus
$\beta(t)\leq 2(1-\nu_{K})^{t}+\left(7+\frac{t}{\pi_{T}(1)}\right)\sum_{x>K}\pi_{T}(x).$
Before we provide the proof, let us discuss various possible regimes:
  • if $\nu_{i}x\rightarrow 0$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ for some $C^{\prime},b>0$ then there is a $C>0$ such that for $x>1$, $\pi(x)=C\exp(-bx)$ and thus
    $\beta(t)$ $\leq 2(1-\nu_{K})^{t}+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}$
    $\leq 2\exp(-t\nu_{K})+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}.$
    For example if $\nu_{x}=\min(1/2,1/\log\log(x))$ we can take $K=t$ to obtain:
    $\beta(t)\leq 2\exp\left(-\frac{t}{\log\log(t)}\right)(1+o(1)).$
  • if $\nu_{x}=\min(1/2,1/x^{a})$ for some $a>0$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ for some $C^{\prime}>0$ and $b>1$ then there is a $C>0$ such that for $x>1$, $\pi_{T}(x)=C\exp(-bx)$ and thus
    $\beta(t)\leq 2\exp\left(-\frac{t}{K^{a}}\right)+\frac{C\left(7+\frac{t}{\pi_{T}(1)}\right)\exp(-b(K+1))}{1-\exp(-b)}$
    which, if we choose $K=\lfloor t^{\frac{1}{a+1}}\rfloor$ we obtain
    $\beta(t)\leq\frac{\frac{Ct}{\pi_{T}(1)}\exp(-bt^{\frac{1}{a+1}})}{1-\exp(-b)}(1+o(1)).$
  • if $\nu_{x}=\min(1/2,1/x)$ and $\mu_{x}=C^{\prime}\nu_{x}/x^{b}$ for some $C^{\prime}>0$ and $b>1$ then there is a $C>0$ such that for $x>1$, $\pi_{T}(x)=C/x^{b+1}$ and thus there is a $C^{\prime\prime}$ with
    $\beta(t)\leq 2\exp(-t/K)+C^{\prime\prime}\left(7+\frac{t}{\pi_{T}(1)}\right)\frac{1}{(K+1)^{b}}$
    and the choice $K=\lfloor t/((b-1)\log t)\rfloor$ gives
    $\beta(t)$ $\leq\left(\frac{2}{t^{d}}+\frac{C^{\prime\prime}(b-1)^{b}}{\pi_{T}(1)}\frac{\log^{b}(t)}{t^{b-1}}\right)(1+o(1))$
    $=\frac{C^{\prime\prime}(b-1)^{b}}{\pi(1)}\frac{\log^{b}(t)}{t^{b-1}}(1+o(1)).$
Of course, these estimates could be seen as pessimistic. However, we also have lower bounds for $\beta(t)$.
Lemma 6.3.
Assume $T=(q,\mu,\nu)\in\mathcal{S}$. Assume $\nu_{x}=\min(1/2,1/x^{a})$ for some $a>0$. Then, we have, for $t\geq 2$,
$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[\frac{1}{4}-\pi_{T}(\lceil t^{1/a}\rceil)\right].$
That is,
$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{8}(1+o(1)).$
This is important with respect to the previous examples:
  • by taking $a=1$ thus $\nu_{x}=\min(1/2,1/x)$ and $\mu(x)=C^{\prime}\nu_{x}/x^{b}=C^{\prime}/x^{b+1}$ for some $b>1$ we have $\pi_{T}(x)=C/x^{b+1}$ and thus
    $\frac{C}{8(t+1)^{b+1}}\left[1+o(1)\right]\leq\beta(t)\leq\frac{C^{\prime\prime}(b-1)^{b}}{\pi_{T}(1)}\frac{\log^{b}(t)}{t^{b-1}}(1+o(1)).$
  • in the case $\nu_{x}=\min(1/2,1/x^{a})$ and $\mu_{x}=C^{\prime}\nu_{x}\exp(-bx)$ and thus $\pi_{T}(x)=C\exp(-bx)$, this gives
    $\frac{C\exp[-b(t^{\frac{1}{a}}+1)]}{8}\left[1+o(1)\right]\leq\beta(t)\leq\frac{\frac{Ct}{\pi_{T}(1)}\exp(-bt^{\frac{1}{a+1}})}{1-\exp(-b)}(1+o(1)).$
We conclude this subsection by the proofs of Lemmas 6.2 and 6.3.
Proof of Lemma 6.2.
The proof relies on the following approximation $P_{T,K}$ of $P_{T}$:
$P_{T,K}=\left(\begin{array}[]{c c c c c c}1-q-q\sum_{x>K+1}\mu_{x}&q\mu_{2}&\dots&q\mu_{K}&0&\dots\\ \nu_{2}&1-\nu_{2}&0&0&0&\dots\\ \vdots&\vdots&\ddots&\vdots&\vdots&\dots\\ \nu_{K}&0&&1-\nu_{K}&0&\dots\\ \nu_{K+1}&0&\dots&0&1-\nu_{K+1}&\dots\\ \vdots&\vdots&&\vdots&\vdots&\ddots\end{array}\right).$
Let $\pi_{T,K}$ denote its invariant distribution:
$\pi_{T,K}(1)=\frac{1}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\text{ and }\pi_{T,K}(x)=\frac{q\mu_{x}}{\nu_{x}\left(1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}\right)}\text{ for }1<x\leq K$
and $\pi_{T,K}(x)=0$ if $x>K$. Then, note that
$\beta(t)$ $=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
$\leq\sum_{x=1}^{K}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}+\sum_{x>K}\pi_{T}(x).$
Then, for $x\leq K$, we have:
$\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}\\ \leq\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}+\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}+\|\pi_{T,K}(\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}.$ (18)
Let us control each term in (18) separately. First,
$2\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}$ $=\left|\frac{1}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}-\frac{1}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\right|$
$\quad\quad\quad+\sum_{x=2}^{K}\left|\frac{\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}-\frac{\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}\right|+\sum_{x>K}\pi_{T}(x)$
$=\left|1-\frac{1+q\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\right|\sum_{x=1}^{K}\pi_{T}(x)+\sum_{x>K}\pi_{T}(x)$
$\leq\left|1-\frac{1+q\sum_{x=2}^{\infty}\frac{\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\right|+\sum_{x>K}\pi_{T}(x)$
$=\frac{\sum_{x>K}\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}+\sum_{x>K}\pi_{T}(x)$
$=\frac{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}\frac{\sum_{x>K}\frac{q\mu_{x}}{\nu_{x}}}{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}+\sum_{x>K}\pi_{T}(x)$
$=\left(\frac{1+q\sum_{y=2}^{\infty}\frac{\mu_{y}}{\nu_{y}}}{1+q\sum_{y=2}^{K}\frac{\mu_{y}}{\nu_{y}}}+1\right)\sum_{x>K}\pi_{T}(x)$
$=\left(\frac{1}{1-\sum_{x>K}\pi_{T}(x)}+1\right)\sum_{x>K}\pi_{T}(x)$
$\leq 3\sum_{x>K}\pi_{T}(x)$
as soon as $K$ large enough to ensure $\sum_{x>K}\pi_{T}(x)\leq 1/2$.
We then tackle the term $\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}$ in (18). In order to do so, we study the Dobrushin coefficient of $P_{T,K}$,
$\kappa:=\sup_{x,y\leq K}\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$
and show it satisfies $\kappa<1$. Exactly as we derived (17) from (16) above, an application of Theorem 18.2.4 in Douc et al. (2018) will allow to conclude. When $1<x<y$,
$\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\left(|\nu_{x}-\nu_{y}|+|1-\nu_{x}|+|1-\nu_{y}|\right)$
$=1-\nu_{y}.$
In the case $1=x<y$, and there we need $q\leq 1/4$ and $\nu_{y}\leq 1/2$,
$\|P_{T,K}(1,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}$ $=\frac{1}{2}\left(\left|1-q-q\sum_{x>K}\mu_{x}-\nu_{y}\right|+|1-\nu_{y}-q\mu_{y}|+\sum_{x\leq K,x\neq y}q\mu_{x}\right)$
$=\frac{1}{2}\left[2-2\nu_{z}-q\left(\sum_{x=2}^{\infty}\mu_{x}+\sum_{x>K}\mu_{x}+\mu_{y}-\sum_{x\leq K,x\neq y}\mu_{x}\right)\right]$
$\leq 1-\nu_{y}.$
So, $\kappa=\sup_{x,y\leq K}\|P_{T,K}(x,\cdot)-P_{T,K}(y,\cdot)\|_{{\rm TV}}=1-\nu_{K}$ which leads to the bound
$\|P_{T,K}^{t}(x,\cdot)-\pi_{T,K}(\cdot)\|_{{\rm TV}}\leq(1-\nu_{K})^{t}.$
Finally, we control the term $\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}$ in (18). This can be done thanks to the theory of the stability of Markov chains, that studies how fast two Markov chains with very similar transition operators will diverge, see for example Mitrophanov (2003). More precisely, (3.17) in the proof of Corollary 3.2 in Mitrophanov (2003) shows that
$\sup_{1\leq x\leq K}\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}\leq t\sup_{1\leq x\leq K}\|P_{T}(x,\cdot)-P_{T,K}(x,\cdot)\|_{{\rm TV}}.$
But it is direct to check that
$\sup_{1\leq x\leq K}\|P_{T}(x,\cdot)-P_{T,K}(x,\cdot)\|_{{\rm TV}}$ $=\|P_{T}(1,\cdot)-P_{T,K}(1,\cdot)\|_{{\rm TV}}$
$=q\sum_{x>K}\mu_{x}.$
Thus, we obtain
$\sup_{1\leq x\leq K}\|P_{T}^{t}(x,\cdot)-P_{T,K}^{t}(x,\cdot)\|_{{\rm TV}}\leq tq\sum_{x>K}\mu_{x}.$
Putting everything together in (18),
$\beta(t)\leq\sum_{x=1}^{K}2\pi_{T}(x)\left[(1-\nu_{K})^{t}+3\sum_{x>K}\pi_{T}(x)+qt\sum_{x>K}\mu_{x}\right]+\sum_{x>K}\pi_{T}(x)$
which ends the proof.
Proof of Lemma 6.3.
We have, for a fixed $t$ and $y$,
$\beta(t)$ $=\sum_{x=1}^{\infty}\pi_{T}(x)\|P_{T}^{t}(x,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
$\geq\pi_{T}(y)\|P_{T}^{t}(y,\cdot)-\pi_{T}(\cdot)\|_{{\rm TV}}$
$=\frac{\pi_{T}(y)}{2}\sum_{y=1}^{\infty}|P_{T}^{t}(k,y)-\pi_{T}(y)|$
$\geq\frac{\pi_{T}(y)}{2}|P_{T}^{t}(y,y)-\pi_{T}(y)|$
$\geq\frac{\pi_{T}(y)}{2}\left[(1-1/(2y^{a}))^{t}-\pi_{T}(y)\right].$
Then, take $y=\lceil t^{1/a}\rceil$ to get
$\beta(t)\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[(1-1/t)^{t}-\pi_{T}(\lceil t^{1/a}\rceil)\right]\geq\frac{\pi_{T}(\lceil t^{1/a}\rceil)}{2}\left[\frac{1}{4}-\pi_{T}(\lceil t^{1/a}\rceil)\right]$
when $t>2$. ∎
6.2   Control of $\mathcal{J}_{p}^{(s)}$ for $P_{T}$
Proposition 6.1.
Assume $T=(q,\mu,\nu)\in\mathcal{T}$ and
$\sum_{x\geq 2}\frac{\mu_{x}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{x}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
Then Condition 3 is satisfied for $P_{T}$ with $D(1)=V(1)=1$, $D(k)=V(k)=1/\sqrt{\mu_{k}}$ for $k>1$ and $b=q\sum_{k\geq 2}\sqrt{\mu_{k}}$. Moreover, the sumability conditions of Lemma 4.1 are also satisfied, and
$\mathcal{J}_{p}^{(s)}\leq\left\|\pi_{T}\right\|_{1-1/p}^{2(1-1/p)}+\left\|1/V\right\|_{1-1/p}^{1-1/p}\left(\left\|\pi_{T}\circ D\right\|_{1-1/p}^{1-1/p}+\left\|\pi_{T}\right\|_{1-1/p}^{1-1/p}[b(s-1)]^{1-1/p}\right)<\infty.$
This gives a nontrivial example of chain that satisfies Condition 3.
Proof of Proposition 6.1.
First, let us evaluate $P_{T}D(x)$. We start with $x>1$,
$P_{T}D(x)=\sum_{y}P_{T}(x,y)D(y)=\nu_{x}D(1)+(1-\nu_{x})D(x)\leq D(x),$
and then when $x=1$,
$P_{T}D(1)=\sum_{y}P_{T}(1,y)D(y)=(1-q)D(1)+q\sum_{y\geq 2}\mu_{y}D(y)\leq D(1)+q\sum_{y\geq 2}\sqrt{\mu_{y}}.$
Thus $P_{T}D(x)\leq D(x)+b$ with $b=q\sum_{y\geq 2}\sqrt{\mu_{y}}$.
We can now check the assumption in Lemma 4.3:
$\sup_{y\in\mathcal{X}}V(y)(P_{T}(x,y)-\pi_{T}(y))\leq D(x).$
We actually prove the stronger:
$\forall(x,y)\text{, }P_{T}(x,y)\leq\frac{D(x)}{V(y)}.$
Indeed, for $x=y=1$,
$P_{T}(1,1)=1-q\leq 1=\frac{D(1)}{V(1)}.$
For $x>1$,
$P_{T}(x,x)=1-\nu_{x}\leq 1=\frac{D(x)}{V(x)}.$
Still for $x>1$,
$P_{T}(x,1)=\nu_{x}\leq 1\leq\frac{1}{\sqrt{\mu_{x}}}=\frac{D(x)}{V(1)}.$
Finally, for $y>1$,
$P_{T}(1,y)=q\mu_{y}\leq\mu_{y}\leq\sqrt{\mu_{y}}=\frac{D(1)}{V(y)}.$
Finally, let us check the summability conditions:
$\sum_{y}\left[\frac{1}{V(y)}\right]^{\frac{1}{2}(1-\frac{1}{p})}=1+\sum_{y>1}\mu_{y}^{\frac{1}{4}(1-\frac{1}{p})}\leq 1+\sum_{y>1}\frac{\mu_{y}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{y}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
by assumption. Then
$\sum_{x}\left[\pi_{T}(x)D_{x}\right]^{\frac{1}{2}(1-\frac{1}{p})}\leq 1+\sum_{x>1}\left[\frac{\mu_{x}}{\nu_{x}}\frac{1}{\sqrt{\mu_{x}}}\right]^{\frac{1}{2}(1-\frac{1}{p})}\leq 1+\sum_{x>1}\frac{\mu_{x}^{\frac{1}{4}(1-\frac{1}{p})}}{\nu_{x}^{\frac{1}{2}(1-\frac{1}{p})}}<\infty$
still by assumption. This ends the proof. ∎
Acknowledgments
GW would like to thank Luca Zanetti and John Sylvester for posing the original question of estimating the average-mixing time, and Aryeh Kontorovich for the interesting discussions.
Funding
GW is supported by the Special Postdoctoral Researcher Program (SPDR) of RIKEN. The visit of Pierre Alquier to RIKEN in October 2023, during which part of this work was conducted was sponsored by the SPDR program.